### v1

# A short note on real algebraic maps which are topologically special generic maps

Abstract.
Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps. A special generic map is, rounghly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
This paper constructs real algebraic maps topologically special generic maps whose images are smoothly embedded manifolds.
We are also interested in construction of explicit and meaningful smooth maps in differential topology and recently ones in real algebraic geometry. This has been an important and difficult problem.
In such stories, we have previously constructed real algebraic maps topologically regarded as special generic maps. This paper is a kind of additional short remark on such maps.
Key words and phrases:
(Non-singular) real algebraic manifolds and real algebraic maps. Smooth maps. Special generic maps.
2020 Mathematics Subject Classification: Primary 14P05, 14P25, 57R45, 58C05. Secondary 57R19.
1. Introduction.
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps. Special generic maps are smooth maps which are locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [4, 7, 27] are related pioneering studies. For theory of Morse functions, see [25] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has also another nice history. [23, 34] show classical and important studies and [1, 2, 16, 17] explain fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
Celebrated theory by Nash, followed by Tognoli, says that smooth closed manifolds are non-singular real algebraic manifolds. They are also shown to be the zero sets of some real polynomial maps. We can also approximate smooth maps between non-singular real algebraic manifolds by real algebraic ones. It is also important and difficult to have real algebraic functions and maps explicitly.
For example, [19, 24, 32] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are seen as simplest symmetric spaces. We are interested in another generalizations of projections of spheres. We are interested in constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [10], followed by [14] by the author himself. We present results of a similar type which are also new.
1.1. Smooth manifolds and maps.
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [22] for example.
Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds. It is also the Riemannian manifold with the standard Euclidean metric. Let $\mathbb{R}:={\mathbb{R}}^{1}$. For $x\in{\mathbb{R}}^{k}$, $||x||\geq 0$ is defined as the canonically defined distance between $x$ and the origin 00. This is also regarded as a (non-singular) real algebraic manifold or the $k$-dimensional real affine space. $S^{k}:=\{x\in{\mathbb{R}}^{k+1}\mid||x||=1\}$ denotes the $k$-dimensional unit sphere, which is a $k$-dimensional smooth compact submanifold of ${\mathbb{R}}^{k+1}$ and has no boundary. It is connected for any positive integer $k\geq 1$. It is a discrete set with exactly two points for $k=0$. It is the zero set of the real polynomial $||x||^{2}={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface). $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk. It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$. Let $S(c)$ denote the set of all singular points of $c$. In our paper, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
Here, a real algebraic manifold means some union of connected components of the intersection of finitely many sets represented as the zero sets of some real polynomial maps. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. Note that we may say Nash manifolds and maps for example for such maps. However, we do not use such terminologies. [29] show related fundamental theory, arguments and some advanced theory.
1.2. Our main results.
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topology. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. See [8] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
Proposition 1 ([27]).
For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
  1. (1)
    We have an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ in such a way that $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
  2. (2)
    We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-N(\partial W_{f})$ whose fiber is the unit sphere.
Main Theorem 1.
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M\rightarrow{\mathbb{R}}^{n})$ of maps on some $m$-dimensional closed and connected manifold $M$ which is also a non-singular real algebraic manifold and the following properties.
  1. (1)
    $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
  2. (2)
    $f_{0}$ is a smooth real algebraic map.
  3. (3)
    There exists a pair $(\Phi,\phi)$ of homeomorphisms enjoying $\Phi\circ f_{0}\circ\phi=f$.
The next section proves Main Theorems. The third section explains about meanings of our main result and related topics by some remarks.
Conflict of interest.
This work was also supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”. The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/). This is closely related to our study. Our study thanks this project for supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He is not employed there. However this helps our studies. Our study also thanks this.
Data availability.
Data essentially supporting our present study are all here.
2. On Main Theorems.
We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00. This is a main theorem of [27].
In [15], followed by [13], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example. Hereafter, for a manifold $X$, ${\rm Int}\ X$ denote the interior $X-\partial X$ of it.
Proposition 2 ([27]).
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given.
  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$
  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
We call the triple of an interior unit sphere bundle for ${\bar{f}}_{N}$, a boundary linear bundle for ${\bar{f}}_{N}$ and this bundle isomorphism a triple for a special generic map and ${\bar{f}}_{N}$ or an SGM-triple for ${\bar{f}}_{N}$. We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have this enjoying the following properties where we can abuse the notation in Proposition 1 and do.
  1. (1)
    The restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ to $\bar{N}-N(\partial\bar{N})\subset\bar{N}-{\rm Int}\ N(\partial\bar{N})$ and the bundle over $\bar{N}-N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
  2. (2)
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ are isomorphic and naturally identified.
  3. (3)
    We also consider the restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in Proposition 1. We consider their natural identification. This coincides with the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$.
Definition 1.
In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle are trivial and the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
[15] has first introduced such a class of special generic maps. We first name the class in this way first in [12]. We do not need to understand this notion or [12, 15] well. However, we note that canonical projections of unit spheres are of such a class. We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers. We choose the connected sum in the smooth category. We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category. This is another simplest example.
We add the following as another main result.
Main Theorem 2.
We can have a pair of maps in Main Theorem 1 as follows.
  1. (1)
    $f$ is product-organized.
  2. (2)
    $M$ is also represented as the zero set of some real polynomial.
A proof of Main Theorems.
We prove our main result. STEPs 1–3 are main ingredients.
STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
This is a new ingredient in our related studies. Our arguments here thanks to [16]. Especially, ”Special Case 5” and ”Definition 14”.
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ by considering a suitable isotopy.
In addition, we can also get the embedding and the subset in such a way that the following are satisfied.
  • ${\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}\geq 0$.
  • ${\rm Int}\ {\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}>0$.
  • At each point of ${\mathbb{R}}^{n}-{\bar{N}}_{0}$, $f_{{\partial}_{j}}<0$ holds for exactly one $1\leq j\leq l$ and $f_{{\partial}_{j}}>0$ holds for the remaining polynomials.
STEP 2 Arguments from [10, 14] to have our smooth real algebraic map $f_{0}$.
We apply arguments from main ingredients of [10, 14].
We consider the zero set $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
In STEP 1, we have seen that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function. Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way. By considering the partial derivative for some variable $x_{i}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We apply the implicit function theorem to see that
$M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
STEP 3 Constructing a special generic map $f$.
For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
We construct a homeomorphism from $M_{0}$ onto another smooth manifold $M$. First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$ where $0\in{\mathbb{R}}^{m-n+1}$ here is the origin of course.
We can consider a smooth function $g:(0,\infty]\rightarrow(0,1]$ enjoying the following.
  • $g(t_{1})\leq g(t_{2})$ for $t_{1}<t_{2}$.
  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$
  • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the standard way.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
We have a homeomorphism onto another manifold $M$. This is smooth and homeomorphic to $M_{0}$. We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
The resulting map is also special generic by the construction and product organized. $M$ is also seen to be ”diffeomorphic” to $M_{0}$. By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems.
This completes the proof. ∎
3. Meanings of our main result and remarks.
The first remark compares our main result to previous results by the author.
Remark 1.
In [10, 11, 14], we consider cases where polynomials are given beforehand. In short, only STEP 2 and STEP 3 are considered. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
The next remark is related to structures of our new real algebraic maps.
Remark 2.
It seems that our new real algebraic maps are not special generic in general in the smooth category. It is important to know whether our maps are special generic or not. It is also difficult.
The following is on another problem.
Remark 3.
As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing canonical projections. Some are closely related to Remark 1.
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. [25] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006 [28] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [18, 20] show explicit answers. [9] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[10] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [3, 30, 31] are also closely related, help the author to do such studies, and motivate us to study further.
4. Acknowledgement.
This new study is based on studies on special generic maps and explicit construction of real algebraic functions and maps of the author. One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry.
The auhtor would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference. He knows singularity theory and applications to geometry widely. Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [12] of the author. He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [33, 35], studies [5, 6] on existence, [26], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [8] for related fundamental and advanced theory from singularity theory for example).
The author would like to thank Atsufumi Honda for asking questions and related informal discussions at Kyushu University after the conference on this theorem, which was an event independent of the previous discussion and held on 2023/12/8. It is also on non-existence of special generic maps on projective spaces and interesting problems from the viewpoint of differential geometry. As a researcher on singularity theory and applications to differential geometry, he says that Reeb’s theorem on Morse functions and special generic maps seem to be closely related to minimizing the total curvatures of curves, surfaces and Riemannian manifolds.
Last, in the informal discussions with Takahiro Yamamoto, the author also says that he is now studying differential geometry and real algebraic geometry further to understand manifolds in related categories, which are finer than the smooth category. We have also discussed the discussions with Atsufumi Honda. He has also commented that special generic maps obtain some nice global structures of manifolds.
In conclusion, we have enjoyed the discussions to present this study.

### v2

@@
-A short note on real algebraic maps which are topologically special generic maps
+A note on real algebraic maps which are topologically special generic maps
 Abstract.
 Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps.
-A special generic map is, rounghly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
+A special generic map is, roughly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
 This paper constructs real algebraic maps topologically special generic maps whose images are smoothly embedded manifolds.
@@
 In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps.
-Such maps are generalized to special generic maps.
-Special generic maps are smooth maps which are locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
+Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
 [4, 7, 28] are related pioneering studies.
-For theory of Morse functions, see [25] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
+For Morse functions, see [26] for a pioneering study and [21] for systematic exposition on fundamental notions and arguments.
 This paper also studies real algebraic geometry.
@@
 For example, [19, 25, 33] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces.
-(Unit) spheres are seen as simplest symmetric spaces.
-We are interested in another generalizations of projections of spheres.
-We are interested in constructing real algebraic maps topologically special generic maps.
+(Unit) spheres are simplest symmetric spaces.
+We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps.
 The author has succeeded in such construction in [10], followed by [14] by the author himself.
@@
 Consult [23] for example.
+Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
 Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds.
@@
 1.2. Our main results.
-Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topology.
+Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies.
 In short two smooth maps are mutually close if their differentials of arbitrary degrees are close.
@@
 Such topologies are important in singularity theory and differential topology.
+We omit rigorous exposition.
 See [8] for systematic exposition.
@@
 A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
-We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
+We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result.
 Proposition 1 ([28]).
@@
 Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists.
-Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M\rightarrow{\mathbb{R}}^{n})$ of maps on some $m$-dimensional closed and connected manifold $M$ which is also a non-singular real algebraic manifold and the following properties.
+Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps on two suitably obtained mutually homeomorphic $m$-dimensional closed and connected manifolds $M$ and $M_{0}$ with the following properties.
   1. (1)
@@
   2. (2)
-    $f_{0}$ is a smooth real algebraic map.
+    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}$.
   3. (3)
-    There exists a pair $(\Phi,\phi)$ of homeomorphisms enjoying $\Phi\circ f_{0}\circ\phi=f$.
+    There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
 The next section proves Main Theorems. The third section explains about meanings of our main result and related topics by some remarks.
@@
 Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given.
+Hereafter we can abuse the notation in Proposition 1 and do.
   • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
-We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$
+We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
   • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
@@
 The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
-We call the triple of an interior unit sphere bundle for ${\bar{f}}_{N}$, a boundary linear bundle for ${\bar{f}}_{N}$ and this bundle isomorphism a triple for a special generic map and ${\bar{f}}_{N}$ or an SGM-triple for ${\bar{f}}_{N}$.
+We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
 We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree.
-Furthermore, we have this enjoying the following properties where we can abuse the notation in Proposition 1 and do.
+Furthermore, we have this enjoying the following properties.
   1. (1)
@@
   3. (3)
-    We also consider the restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in Proposition 1.
+    We also consider the restriction of bundle ${\bar{M}}_{\rm I}$ in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here.
 We consider their natural identification.
-This coincides with the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$.
+This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
 Definition 1.
-In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle are trivial and the bundle isomorphism in the SGM-triple for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
+In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
 [15] has first introduced such a class of special generic maps.
 We first name the class in this way first in [12].
-We do not need to understand this notion or [12, 15] well.
+Of course in our present paper, we do not need to understand this notion or [12, 15] well and we do not assume their arguments or results.
 However, we note that canonical projections of unit spheres are of such a class.
@@
 We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category.
-This is another simplest example.
-We add the following as another main result.
+This gives another simplest example.
+We add the following as additional main results.
 Main Theorem 2.
@@
   2. (2)
-    $M$ is also represented as the zero set of some real polynomial.
+    $M_{0}$ is also represented as the zero set of some real polynomial.
+Main Theorem 3.
+In Main Theorem 2, we can have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
+  1. (1)
+    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” and have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and whose structure group is reduced to a group consisting of linear transformations of $D^{m-n+1}$: in fact a manifold homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined except for $m-n+1=k\neq 4$.
+  2. (2)
+    Under the structure group, the bundle of (1) here is in fact trivial.
+  3. (3)
+    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ diffeomorphic to a manifold obtained in the following steps.
+    1. (a)
+      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ homeomorphic to $D^{m-n+2}$.
+Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
+    2. (b)
+      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
+    3. (c)
+      Eliminate the corner.
+    In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
+Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered.
 A proof of Main Theorems.
-We prove our main result. STEPs 1–3 are main ingredients.
+We prove our main result. First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
 STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
@@
 is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$.
-The map $f_{0}$ is obtained as the composition of the canonical embedding of $M\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
+The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
 STEP 3 Constructing a special generic map $f$.
@@
 The resulting map is also special generic by the construction and product organized.
-$M$ is also seen to be ”diffeomorphic” to $M_{0}$.
-By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems.
+By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
+This completes the proof of Main Theorems 1 and 2.
+We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
+We first show Main Theorem 3 (1).
+First the bundle we discuss here is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
+We apply an argument used to show Proposition 1.
+More precisely, Proposition 2.1 of [28] is a corresponding result.
+It also refers to [22] there.
+The structure group is regarded as a group consisting of diffeomorphisms preserving the distance between each point and the origin $0\in{\Delta}^{m-n+1}$ where we can define the origin 00 for ${\Delta}^{m-n+1}$ not diffeomorphic to $D^{m-n+1}$.
+We explain about this.
+For an element ${\phi}_{{\Delta}^{m-n+1}}$ of such a group, we consider $\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)$ for $x\in{\Delta}^{m-n+1}$ and $0<t\leq 1$ where $x$ and $tx$ are canonically regarded as vectors.
+Here $||\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)||=||x||$.
+We consider the limit at $t=0$.
+For a fixed $x$, we have the value of the differential of ${\phi}_{{\Delta}^{m-n+1}}$ at $x$, seen naturally as a tangent vector at the origin of ${\Delta}^{m-n+1}$.
+Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”.
+This means that the structure group reduces to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
+This completes the proof of Main Theorem 3 (1).
+We see that this bundle is trivial for the structure group.
+$W_{0}$ is defined as
+$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
+and this can be also defined as
+$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$.
+We use the function $g$ before.
+We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule where we consider some elements as vectors.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
+  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
+From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$.
+By the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$, we can see that the previous bundle is trivial for the structure group.
+This completes the proof of Main Theorem 3 (2).
+By our arguments, we can also check Main Theorem 3 (3) easily.
+A result close to this is also presented in Theorem 1 (3) of [14].
+”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there.
+The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there: only exposition on ”homotopy” is presented there.
 This completes the proof. ∎
@@
 In [10, 11, 14], we consider cases where polynomials are given beforehand.
-In short, only STEP 2 and STEP 3 are considered.
+In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark and this does not affect main ingredients of [10, 11, 14].
+STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first.
 These studies also consider explicit cases.
@@
 For this, [3, 31, 32] are also closely related, help the author to do such studies, and motivate us to study further.
+The following is a remark on Main Theorem 3.
+Remark 4.
+As another method, we can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively.
+In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
 4. Acknowledgement.
@@
 One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry.
-The auhtor would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference.
+The author would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference.
 He knows singularity theory and applications to geometry widely.

### v3

@@
 Real algebraic geometry has also another nice history.
-[24, 35] show classical and important studies and [1, 2, 16, 17] explain fundamental notions and arguments and advanced studies, for example.
+[34, 46] show classical and important studies and [1, 2, 17, 18] present fundamental notions and arguments and advanced studies, for example.
 Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
@@
 Consult [33] for example.
-Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
+Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ denote the interior of it.
 Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds.
@@
 $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$.
-$c(x)$ is called a singular value of $c$.
+$c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$.
 Let $S(c)$ denote the set of all singular points of $c$.
-In our paper, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth.
+Here, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth.
 A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$.
@@
 [41] show related fundamental theory, arguments and some advanced theory.
-1.2. Our main results.
+1.2. Our main result and the structure of our paper.
 Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies.
@@
 A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group.
-A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
-We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result.
+A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
+We present a theorem on fundamental structures of special generic maps.
+Using the notation, we also present our main result as Main Theorem 1.
+This is proven in the next section with additional results Main Theorems 2 and 3.
+The third section is on meanings of our main result and related topics.
 Proposition 1 ([39]).
@@
     We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$.
-By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-N(\partial W_{f})$ whose fiber is the unit sphere.
+By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
 Main Theorem 1.
@@
   2. (2)
-    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}$.
+    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$.
   3. (3)
     There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
-The next section proves Main Theorems. The third section explains about meanings of our main result and related topics by some remarks.
 Conflict of interest.
@@
 In [16], followed by [14], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example.
-Hereafter, for a manifold $X$, ${\rm Int}\ X$ denote the interior $X-\partial X$ of it.
 Proposition 2 ([39]).
@@
 More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$.
-Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$.
+Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as $\partial\bar{N}\times\{1\}$ naturally and the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$.
 The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
@@
   1. (1)
-    The restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ to $\bar{N}-N(\partial\bar{N})\subset\bar{N}-{\rm Int}\ N(\partial\bar{N})$ and the bundle over $\bar{N}-N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
+    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
   2. (2)
@@
   3. (3)
-    We also consider the restriction of bundle ${\bar{M}}_{\rm I}$ in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here.
+    We also consider the restriction of the smooth bundle in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here.
 We consider their natural identification.
@@
 We first name the class in this way first in [13].
-Of course in our present paper, we do not need to understand this notion or [12, 15] well and we do not assume their arguments or results.
+Of course in our present paper, we do not need to understand this notion or [13, 16] well.
+We do not assume their arguments or results.
 However, we note that canonical projections of unit spheres are of such a class.
@@
   2. (2)
-    $M_{0}$ is also represented as the zero set of some real polynomial.
+    $M_{0}\subset{\mathbb{R}}^{m+1}$ is also represented as the zero set of some real polynomial.
 Main Theorem 3.
@@
   1. (1)
-    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” and have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and whose structure group is reduced to a group consisting of linear transformations of $D^{m-n+1}$: in fact a manifold homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined except for $m-n+1=k\neq 4$.
+    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
+We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, the boundary of whose fiber is diffeomorphic to $S^{m-n}$ and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$: in fact a smooth manifold which is homeomorphic to $D^{k}$ and whose fiber is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined before except for $m-n+1=k\neq 4$.
   2. (2)
-    Under the structure group, the bundle of (1) here is in fact trivial.
+    Under the presented structure group, the bundle of (1) here is in fact trivial.
   3. (3)
@@
     1. (a)
-      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ homeomorphic to $D^{m-n+2}$.
+      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is diffeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$.
 Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
@@
     In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
-Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered.
+Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$.
 A proof of Main Theorems.
-We prove our main result. First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
+First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
 STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
@@
 By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [19].
+For this approximation, see also [3, 7].
 This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$.
-Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$.
+Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$.
 Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$.
-We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ by considering a suitable isotopy.
+We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
 In addition, we can also get the embedding and the subset in such a way that the following are satisfied.
@@
 We consider the zero set $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
-In STEP 1, we have seen that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
-Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function.
+The last part of STEP 1 implies that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
+Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
 Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
-By considering the partial derivative for some variable $x_{i}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value.
+By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value.
+We explain about this more precisely.
+If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$.
+This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials.
+Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$.
+The first value is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class.
+The second value is not zero by the property of the zero sets of the polynomials.
 We apply the implicit function theorem to see that
 $M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
-is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$.
+is a smooth, closed and connected manifold and non-singular real algebraic manifold of dimension $m$.
+Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$.
 The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
   • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
-We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the standard way.
+We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the canonical way.
   • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
@@
 We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
-The resulting map is also special generic by the construction and product organized.
-By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
+The resulting map is also special generic by the construction.
+It is also product-organized.
+By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
 This completes the proof of Main Theorems 1 and 2.
 We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
-We first show Main Theorem 3 (1).
-First the bundle we discuss here is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
-We apply an argument used to show Proposition 1.
-More precisely, Proposition 2.1 of [28] is a corresponding result.
-It also refers to [22] there.
-The structure group is regarded as a group consisting of diffeomorphisms preserving the distance between each point and the origin $0\in{\Delta}^{m-n+1}$ where we can define the origin 00 for ${\Delta}^{m-n+1}$ not diffeomorphic to $D^{m-n+1}$.
-We explain about this.
-For an element ${\phi}_{{\Delta}^{m-n+1}}$ of such a group, we consider $\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)$ for $x\in{\Delta}^{m-n+1}$ and $0<t\leq 1$ where $x$ and $tx$ are canonically regarded as vectors.
-Here $||\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)||=||x||$.
-We consider the limit at $t=0$.
-For a fixed $x$, we have the value of the differential of ${\phi}_{{\Delta}^{m-n+1}}$ at $x$, seen naturally as a tangent vector at the origin of ${\Delta}^{m-n+1}$.
-Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”.
-This means that the structure group reduces to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
-This completes the proof of Main Theorem 3 (1).
-We see that this bundle is trivial for the structure group.
-$W_{0}$ is defined as
+We define $W_{0}$ as
 $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
@@
 From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$.
-By the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$, we can see that the previous bundle is trivial for the structure group.
+We investigate the bundle in Main Theorem 3 (1) and (2).
+Let the bundle denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$.
+First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
+We apply an argument used to show Proposition 1 in a revised style.
+More precisely, Proposition 2.1 of [39] is a corresponding result.
+It also refers to [32] there.
+We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$.
+We remove the total space of this bundle obtained by the restriction from $W_{0}$.
+We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$.
+This is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$.
+This is also thanks to Ehresmann’s fibration theorem (or a relative version of this).
+Hereafter we apply this implicitly in several scenes.
+The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$.
+The structure group of the bundle $M_{{\rm B},0}$ is restricted to the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$.
+For each $0<t\leq 1$, it is a subbundle of the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly.
+For each two obtained smooth bundles, fibers are regarded as $\partial D^{m-n+1}$ and $D^{m-n+1}$, respectively.
+The former bundle is also obtained by restricting the fiber to the boundary $\partial D^{m-n+1}$ of course.
+We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before.
+This element can be regarded as an element extended to an element preserving the distance between each point and the origin $0\in D^{m-n+1}$.
+This is also regarded as an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
+For such an element, we can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}$ which is not the origin and $0<s\leq 1$.
+Of course $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors and hereafter we do not explain about this rule explicitly.
+Here $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$.
+We consider the limit at $s=0$.
+For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$.
+Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”.
+This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
+These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1).
+We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group.
 This completes the proof of Main Theorem 3 (2).
@@
 In [11, 12, 15], we consider cases where polynomials are given beforehand.
-In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark and this does not affect main ingredients of [10, 11, 14].
+In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark.
+This does not affect main ingredients of [11, 12, 15].
 STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first.
@@
 It seems that our new real algebraic maps are not special generic in general in the smooth category. It is important to know whether our maps are special generic or not. It is also difficult.
+As another problem which is in some sense similar to this problem, it is also very difficult to know whether two homeomorphisms $\Phi$ and $\phi$ in Main Theorem 1 can be chosen as diffeomorphisms in general.
 The following is on another problem.
@@
 The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages.
+The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
 [36] is a pioneering study considering Reeb graphs.
 Such tools are fundamental and strong tools in studying manifolds by using graphs.
-In 2006 [29] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph.
+In 2006, [40] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph.
 [20, 30] show explicit answers.
@@
 For this, [3, 42, 43] are also closely related, help the author to do such studies, and motivate us to study further.
+They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
+Related to Remark 3, we add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [22, 23, 24, 25, 26, 27, 28, 29].
+See also [37].
+This is also closely related to [3].
+This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory and this mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs.
+We do not explain about elementary notions on graphs precisely.
+We except related knowledge.
+We also note that we omit several rigorous discussions in Remark 4 and that for our main ingredients of the present paper we do not need to understand this rigorously.
+Remark 4.
+[3] considers a finite graph $K$ enjoying the following properties.
+  • We can embed $K$ by a piecewise smooth embedding $e:K\rightarrow{\mathbb{R}}^{2}$.
+  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is an smooth embedding.
+  • The degree of each vertex is $1$ or $3$.
+  • At the vertex where the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ has a local extremum, the degree of the vertex is $1$.
+An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves.
+The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of connected components of all preimages of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$ under the natural equivalence relation.
+Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$.
+The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ containing $e(K)$ and its Poincaré-Reeb graph is isomorphic to and naturally regarded as $e(K)$.
+Moreover, we can have the situation in the following.
+  1. (1)
+    The restriction of ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
+  2. (2)
+    $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
+This is also obtained by approximation of the Weierstrass type in the class $C^{2}$.
+According to arguments similar to one in ”Special Case 5” of [17], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
 The following is a remark on Main Theorem 3.
@@
 In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
+We give a short remark on [15] last.
+Remark 6.
+[15] has no essential errors and we have also found several more rigorous arguments.
+These arguments are discussed in the present paper.
+We will revise the version and submit the refereed journal based on comments and our arguments and idea here.
+Note also that the content of [15] is independent of the content of the present paper essentially.
 4. Acknowledgement.

### v4

@@
-A note on real algebraic maps which are topologically special generic maps
+Some remarks on real algebraic maps which are topologically special generic maps
 Abstract.
@@
 We are also interested in construction of explicit and meaningful smooth maps in differential topology and recently ones in real algebraic geometry. This has been an important and difficult problem.
-In such stories, we have previously constructed real algebraic maps topologically regarded as special generic maps. This paper is a kind of additional short remark on such maps.
+In such stories, we have previously constructed some real algebraic maps topologically regarded as special generic maps.
+This paper constructs such maps in general situations and is mainly for additional remarks on such maps.
 Key words and phrases:
@@
 The author has succeeded in such construction in [14], followed by [17] by the author himself.
-We present results of a similar type which are also new.
+We present new related results: related new examples.
 1.1. Smooth manifolds and maps.
@@
 A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
-Here, a real algebraic manifold means some union of connected components of the intersection of finitely many sets represented as the zero sets of some real polynomial maps.
+Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map.
 Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps.
@@
 This is proven in the next section with additional results Main Theorems 2 and 3.
+Main Theorem 3 is a more generalized result.
 The third section is on meanings of our main result and related topics.
 Proposition 1 ([43]).
+Let $m>n\geq 1$ be integers.
 For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
@@
   2. (2)
-    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$.
+    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ which is also the zero set of a real polynomial.
   3. (3)
-    There exists a pair $(\Phi;M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$.
+    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$.
+This paper is regarded as an improved version of [18] and based on our new discovery and idea.
 Conflict of interest.
@@
 Hereafter we can abuse the notation in Proposition 1 and do.
+We can also abuse this in the case $m=n$ and we do.
   • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
@@
   2. (2)
-    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ are isomorphic and naturally identified.
+    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified.
   3. (3)
@@
 This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
+Note that in Proposition 1, we cannot set ”$m\geq n$”. This is closely related to Eliashberg’s celebrated theory [6]. See also [7].
 Definition 1.
@@
 This gives another simplest example.
-We add the following as additional main results.
+We add the following as another additional main result.
 Main Theorem 2.
-We can have a pair of maps in Main Theorem 1 as follows.
-  1. (1)
-    $f$ is product-organized.
-  2. (2)
-    $M_{0}\subset{\mathbb{R}}^{m+1}$ is also represented as the zero set of some real polynomial.
-Main Theorem 3.
-In Main Theorem 2, we can have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
-  1. (1)
-    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
-We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, the boundary of whose fiber is diffeomorphic to $S^{m-n}$ and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$: in fact a smooth manifold which is homeomorphic to $D^{k}$ and whose fiber is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined before except for $m-n+1=k\neq 4$.
-  2. (2)
-    Under the presented structure group, the bundle of (1) here is in fact trivial.
-  3. (3)
-    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ diffeomorphic to a manifold obtained in the following steps.
-    1. (a)
-      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is diffeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$.
-Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
-    2. (b)
-      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
-    3. (c)
-      Eliminate the corner.
-    In fact, except the case where ${\Delta}^{k}$ is not diffeomorphic to $D^{k}$, $M$ and $M_{0}$ are diffeomorphic.
-Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$.
-A proof of Main Theorems.
-First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
+In Main Theorem 1, $f$ and $f_{0}$ can be obtained as product-organized special generic maps.
+A proof of Main Theorems 1 and 2.
+STEPs 1–3 are main ingredients.
 STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
@@
 Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$.
-The first value is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class.
-The second value is not zero by the property of the zero sets of the polynomials.
+The value of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ for $x_{j_{0}}$ is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class.
+The product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$ is not zero by the property of the zero sets of the polynomials.
 We apply the implicit function theorem to see that
 $M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
-is a smooth, closed and connected manifold and non-singular real algebraic manifold of dimension $m$.
+is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
 Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$.
 The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
-STEP 3 Constructing a special generic map $f$.
+STEP 3 Is $f_{0}$ special generic?
+We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $\partial{\bar{N}}_{0}\times[0,1]\subset N(\partial{\bar{N}}_{0})$.
+More precisely, we can have one and consider identification in the following way thanks to arguments related to ”Definition 14” [20].
+  • $\partial{\bar{N}}_{0}\times\{0\}$ is identified with $\partial N(\partial{\bar{N}}_{0})\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
+  • $\partial{\bar{N}}_{0}\times\{t_{0}\}\subset\partial{\bar{N}}_{0}\times(0,1]$ is identified with the zero set of the polynomial $f_{j}-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$.
+Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function: some $y_{j}$ is regarded as the variable of the target of the function.
+We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin.
+This is a diffeomorphism.
+We map each point of the form $(x,0)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
+Thus we have a diffeomorphism from ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ onto ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ and by the restriction another diffeomorphism $({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto ${\rm Int}\ {\bar{N}}_{0}\times S^{m-n}\subset{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
+These arguments imply that $f_{0}$ is a product-organized special generic map.
+We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3.
+By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
+This completes the proof.
+Main Theorem 3.
+Suppose a situation as in Main Theorem 1 is given.
+Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following.
+We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $t$.
+  • $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
+  • $F_{0,S^{m-n}}(0)=0$. $F_{0,S^{m-n}}(t)>0$ for $t>0$ and $F_{0,S^{m-n}}(t)<0$ for $t<0$. The value of the derivative of $F_{0,S^{m-n}}$ at 00 is not zero.
+  • ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$. $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)\geq 0$ for $y\in{\mathbb{R}}^{m-n+1}$.
+  • ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the empty set for $t<0$.
+  • For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$.
+We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)$ mapping ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively.
+${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
+  • For any real number $\epsilon>0$, we have a positive real number $\delta>0$ and for $0<t<\delta$ and each point $p$ of ${F_{S^{m-n},t}}^{-1}(0)$, $0<||p||<\epsilon$.
+Then we can have a statement of Main Theorem 1 by defining
+$M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
+for a sufficiently large $T>0$.
+As a more specific case, we can also have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
+  1. (1)
+    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
+We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$.
+This is also a bundle which may not be smooth and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$.
+  2. (2)
+    Under the presented structure group, the bundle of (1) here which may not be a smooth bundle is in fact trivial.
+  3. (3)
+    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps.
+    1. (a)
+      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is homeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$.
+Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4,5$.
+Note also that for $k\neq 4$ this is diffeomorphic to $D^{k}$ if its boundary is diffeomorphic to $S^{k-1}$.
+    2. (b)
+      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
+    3. (c)
+      Eliminate the corner.
+We review several known facts on topologies and differentiable structures of spheres and disks shortly.
+See [34] and see also [12, 46, 47] for example.
+A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example.
+It is well-known that smooth manifolds homeomorphic to $S^{k}$ are classified via abstract and sophisticated theory for $k\geq 7$.
+If a smooth manifold which is homeomorphic to $S^{k}$ is smoothly embedded into ${\mathbb{R}}^{k+1}$ for $k\geq 1$ with $k\neq 4$, then this is diffeomorphic to $S^{k}$.
+A smooth manifold ”which is homeomorphic to $S^{4}$ and not diffeomorphic to $S^{4}$” is still undiscovered.
+A smooth manifold which is homeomorphic to $D^{k}$ and whose boundary is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ unless $k=4$.
+A smooth manifold which is homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ unless $k=4,5$.
+Remember that ”${\Delta}^{k}$ which is homeomorphic to $D^{k}$ and is not diffeomorphic to $D^{k}$” is still undiscovered.
+If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
+A proof of Main Theorem 3.
+We can argue as STEP 1 of the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
+We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than $a$. We choose another positive number $T_{0}<a$.
+We present an argument similar to one of STEP 2 of the proof of Main Theorems 1 and 2.
+Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
+This mainly comes from the following assumption: ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
+Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 of the proof of Main Theorems 1 and 2.
+The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not zero from the assumption, and the value in the case of Main Theorems 1 and 2.
+The resulting value is not zero.
+We can argue similarly to prove that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
+We present an argument similar to one of STEP 3 of the proof of Main Theorems 1 and 2.
 For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
-We construct a homeomorphism from $M_{0}$ onto another smooth manifold $M$.
-First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$ where $0\in{\mathbb{R}}^{m-n+1}$ here is the origin of course.
-We can consider a smooth function $g:(0,\infty]\rightarrow(0,1]$ enjoying the following.
+We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
+First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$.
+Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
+We can consider a smooth function $g:(0,\infty)\rightarrow(0,1]$ enjoying the following.
   • $g(t_{1})\leq g(t_{2})$ for $t_{1}<t_{2}$.
-  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$
+  • $g(t)=\sqrt{1-{(1-t)}^{2}}$ for a sufficiently small $t>0$.
   • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
-We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the canonical way.
-  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
+We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
+$(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$
+according to the following rule.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
   • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
-We have a homeomorphism onto another manifold $M$.
-This is smooth and homeomorphic to $M_{0}$.
+We have a homeomorphism onto a smooth manifold $M$.
 We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
 It is also product-organized.
-By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
-This completes the proof of Main Theorems 1 and 2.
-We show Main Theorem 3. We can prove this by applying (revised versions of) known arguments previously presented in related studies and our present study.
 We define $W_{0}$ as
-$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$
+$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
 and this can be also defined as
-$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}\geq 0\}$.
+$\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$.
 We use the function $g$ before.
-We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule where we consider some elements as vectors.
+We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
+This is a diffeomorphism.
+We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
+We thus have a homeomorphism from $W_{0}$.
+The image of the homeomorphism is the union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
   • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
   • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
-From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$.
+From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
+We can define another non-singular real algebraic manifold
+$M_{0,t}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t)=0\}$
+for $0<t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. This is diffeomorphic to $M_{0}$. This new set has properties similar to important ones of $M_{0}$.
+We go back to the trivial bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle.
+This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
+In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, this smooth bundle which may not be trivial in general is also trivial.
+Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
 We investigate the bundle in Main Theorem 3 (1) and (2).
-Let the bundle denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$.
+We can present a simple and natural proof.
+However, we present a complicated proof to respect general cases.
+Let the bundle in Main Theorem 3 (1) and (2) denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$.
 First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
@@
 We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$.
-We remove the total space of this bundle obtained by the restriction from $W_{0}$.
+From $W_{0}$, we remove the total space of this bundle over ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$.
 We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$.
-This is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$.
+This fiber is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$.
 This is also thanks to Ehresmann’s fibration theorem (or a relative version of this).
-Hereafter we apply this implicitly in several scenes.
 The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$.
-The structure group of the bundle $M_{{\rm B},0}$ is restricted to the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$.
-For each $0<t\leq 1$, it is a subbundle of the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly.
-For each two obtained smooth bundles, fibers are regarded as $\partial D^{m-n+1}$ and $D^{m-n+1}$, respectively.
-The former bundle is also obtained by restricting the fiber to the boundary $\partial D^{m-n+1}$ of course.
+The structure group of the bundle $M_{{\rm B},0}$ is naturally regarded as the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$.
+For each $0<t\leq 1$, it is a subbundle of the following bundle: the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly.
+For each two obtained smooth bundles for $0<t\leq 1$, fibers are regarded as a manifold homeomorphic to $\partial D^{m-n+1}$ and a manifold homeomorphic to $D^{m-n+1}$, respectively.
+The former bundle is also obtained by restricting the fiber to the boundary homeomorphic $\partial D^{m-n+1}$ of course.
+The manifold homeomorphic to $D^{m-n+1}$ is diffeomorphic to $D^{m-n+1}$ if $m-n+1\neq 4,5$ or the boundary is diffeomorphic to $\partial D^{m-n+1}=S^{m-n}$ with the constraint $m-n+1\neq 4$ of course.
 We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before.
-This element can be regarded as an element extended to an element preserving the distance between each point and the origin $0\in D^{m-n+1}$.
-This is also regarded as an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
-For such an element, we can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}$ which is not the origin and $0<s\leq 1$.
-Of course $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors and hereafter we do not explain about this rule explicitly.
-Here $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$.
+This element can be regarded as an element canonically extended to an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
+By these arguments, we may discuss the bundles and the structure groups in the following simple situation essentially.
+We can regard the structure group of the bundle $W_{1,t}$ as a group acting on the unit disk $D^{m-n+1}$ preserving the distance of each point and the origin $0\in D^{m-n+1}$.
+We explain about an element ${\psi}_{t}$ of the group.
+We can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}-\{0\}$ and $0<s\leq 1$.
+Of course 00 here is the origin and $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors.
+Here we may regard the relation $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$ by considering the distance $||x||$ for $x\in D^{m-n+1}-\{0\}\subset{\mathbb{R}}^{m-n+1}$.
 We consider the limit at $s=0$.
@@
 ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there.
-The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there: only exposition on ”homotopy” is presented there.
+The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there.
+Only exposition on ”homotopy” is presented there.
 This completes the proof. ∎
+Example 1.
+Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$.
+For this we can choose $a$ as an arbitrary positive real number.
+The bundles in (1) and (2) are regarded as ”linear bundles defined before”.
+Furthermore, in this case, we can consider a suitable situation and regard that ${\Phi}_{F_{S^{m-n}},a}$ is replaced by a diffeomorphism, that the homeomorphism mapping each sphere is regarded as a diffeomorphism, and that (3) can be discussed in the smooth category, for example.
+In short, we can argue all in the smooth situation in Main Theorem 3.
+This is also for Main Theorem 2.
+This is generalized for more general cases with $m-n\neq 4$ for example.
+This is thanks to fundamental arguments on singularity theory and differential topology.
+This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres.
+This is also related to studies on diffeomorphism groups of unit spheres, studied in [11], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
+See also Remark 4, in the next section, for example.
+Example 2.
+Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a Morse function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$.
+We have a special generic map $f_{0}$ as in the case of Main Theorem 2.
+We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with $k_{j,1}$ being positive numbers and $k_{j,2}$ being positive integers.
+See also Remark 3 of [14].
 3. Meanings of our main result and remarks.
@@
 These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
-The next remark is related to structures of our new real algebraic maps.
-Remark 2.
-It seems that our new real algebraic maps are not special generic in general in the smooth category. It is important to know whether our maps are special generic or not. It is also difficult.
-As another problem which is in some sense similar to this problem, it is also very difficult to know whether two homeomorphisms $\Phi$ and $\phi$ in Main Theorem 1 can be chosen as diffeomorphisms in general.
 The following is on another problem.
@@
 [17] has no essential errors and we have also found several more rigorous arguments.
-These arguments are discussed in the present paper.
-We will revise the version and submit the refereed journal based on comments and our arguments and idea here.
+Some of these important arguments are discussed in the present paper.
+We will revise the version and submit it to the refereed journal, based on comments and our arguments and idea here.
 Note also that the content of [17] is independent of the content of the present paper essentially.
@@
 One of related events motivating the author to present this study is the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)” and related discussions on singularity theory, differential geometry, and real algebraic geometry.
-The author would like to thank Takahiro Yamamoto for asking some questions on the talk by the author, given on 2023/11/27, and related informal discussions at Kyushu University held on 2023/12/11, after the conference.
-He knows singularity theory and applications to geometry widely.
-Based on his knowledge and idea, he has first commented that projective spaces admit no special generic maps in considerable cases and that it is interesting: this is on the study [13] of the author.
-He also says that it is also important to consider so-called general fold maps and explicit classes for understanding projective spaces in geometric ways: fold maps are defined naturally generalizing the form of the last ($n$-th) component in the definition of a special generic map (see related pioneering studies [45, 47], studies [5, 6] on existence, [38], a pioneering studies on algebraic topology and differential topology of manifolds admitting fold maps, and [9] for related fundamental and advanced theory from singularity theory for example).
-The author would like to thank Atsufumi Honda for asking questions and related informal discussions at Kyushu University after the conference on this theorem, which was an event independent of the previous discussion and held on 2023/12/8.
-It is also on non-existence of special generic maps on projective spaces and interesting problems from the viewpoint of differential geometry.
-As a researcher on singularity theory and applications to differential geometry, he says that Reeb’s theorem on Morse functions and special generic maps seem to be closely related to minimizing the total curvatures of curves, surfaces and Riemannian manifolds.
-Last, in the informal discussions with Takahiro Yamamoto, the author also says that he is now studying differential geometry and real algebraic geometry further to understand manifolds in related categories, which are finer than the smooth category.
-We have also discussed the discussions with Atsufumi Honda.
-He has also commented that special generic maps obtain some nice global structures of manifolds.
-In conclusion, we have enjoyed the discussions to present this study.
+See also the previous versions [18].

### v5

@@
-Some remarks on real algebraic maps which are topologically special generic maps
+Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres
 Abstract.
-Morse functions with exactly two singular points on homotopy spheres and canonical projections of spheres are generalized as special generic maps.
-A special generic map is, roughly, a smooth map represented as the composition of a smooth surjection onto a manifold whose preimages are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary with a smooth immersion of codimension 00.
-This paper constructs real algebraic maps topologically special generic maps whose images are smoothly embedded manifolds.
-We are also interested in construction of explicit and meaningful smooth maps in differential topology and recently ones in real algebraic geometry. This has been an important and difficult problem.
-In such stories, we have previously constructed some real algebraic maps topologically regarded as special generic maps.
-This paper constructs such maps in general situations and is mainly for additional remarks on such maps.
+Morse functions with exactly two singular points on spheres and canonical projections of spheres are generalized to special generic maps.
+A special generic map is, roughly, a smooth map represented as the composition of a surjection onto a manifold with an immersion of codimension 00.
+Preimages of this surjections are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary.
+We construct real algebraic maps topologically seen as special generic maps whose images are smoothly embedded manifolds.
+Canonical projections of unit spheres are generalized here.
+We are interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry.
+Obtaining them explicitly is important and difficult.
+In the stories, we previously constructed some real algebraic maps topologically special generic maps.
+We construct such maps in more general situations and give additional remarks.
 Key words and phrases:
-(Non-singular) real algebraic manifolds and real algebraic maps. Smooth maps. Special generic maps.
+(Non-singular) real algebraic manifolds and real algebraic maps. Real algebraic hypersurfaces. Hilbert’s 16th problem. Smooth maps. Special generic maps.
 2020 Mathematics Subject Classification: Primary 14P05, 14P25, 57R45, 58C05. Secondary 57R19.
@@
 In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps.
-Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or singular points of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
+Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
 [5, 9, 44] are related pioneering studies.
 For Morse functions, see [41] for a pioneering study and [36] for systematic exposition on fundamental notions and arguments.
+For higher dimensional versions of theory of Morse functions, [52, 55] are pionnering studies.
 This paper also studies real algebraic geometry.
 Complex algebraic geometry is presenting one of central topics in algebraic geometry.
-Real algebraic geometry has also another nice history.
+Real algebraic geometry has another nice history.
 [39, 53] show classical and important studies and [2, 3, 21, 22] present fundamental notions and arguments and advanced studies, for example.
-Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
+Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of) spheres.
 Celebrated theory by Nash, followed by Tognoli, says that smooth closed manifolds are non-singular real algebraic manifolds.
@@
 It is also important and difficult to have real algebraic functions and maps explicitly.
-For example, [24, 39, 50] shows explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces.
+For example, [25, 40, 51] show explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces.
 (Unit) spheres are simplest symmetric spaces.
-We are interested in another generalizations of projections of spheres: constructing real algebraic maps topologically special generic maps.
+We consider another generalization for projections of spheres: constructing real algebraic maps topologically special generic maps.
 The author has succeeded in such construction in [15], followed by [18] by the author himself.
-We present new related results: related new examples.
+We present new related result.
+More precisely, we construct new explicit maps on some real algebraic hypersurfaces.
+As one of related future work, [10, 54] present closely related long-standing studies on construction of real algebraic curves in the plane and general real algebraic hypersurfaces (Remark 5).
 1.1. Smooth manifolds and maps.
 Let $X$ be a topological space with the structure of a cell complex whose dimension is finite.
-We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer.
+We can define the dimension $\dim X$ uniquely, which is a non-negative integer.
 A topological manifold is well-known to be homeomorphic to some CW complex.
@@
 Consult [38] for example.
-Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ denote the interior of it.
+Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ the interior of it.
 Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds.
@@
 It is a discrete set with exactly two points for $k=0$.
-It is the zero set of the real polynomial $||x||^{2}={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface).
+It is the zero set of the real polynomial $||x||^{2}-1={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}-1$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface).
 $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk.
 It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
-Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$.
+Let $c:X\rightarrow Y$ be a differentiable map.
 $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$.
@@
 This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$.
-A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
+The canonical projection of the unit sphere $S^{k-1}$ is its restriction.
 Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map.
@@
 Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections.
-Note that we may say Nash manifolds and maps for example for such maps.
-However, we do not use such terminologies.
-[45] show related fundamental theory, arguments and some advanced theory.
+We may say Nash manifolds and maps for example for such maps.
+[46] shows related fundamental theory, arguments and some advanced theory.
 1.2. Our main result and the structure of our paper.
-Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies.
-In short two smooth maps are mutually close if their differentials of arbitrary degrees are close.
-$C^{r}$ Whitney topologies are also considered as weaker topologies.
-They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course.
+Spaces of smooth maps between two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies.
+In short two smooth maps are close if their differentials of arbitrary degrees are close.
+$C^{r}$ Whitney topologies are also considered as weaker topologies for each non-negative integer $r$.
+They respect differentials of degrees at most $r$.
 The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$.
@@
 See [11] for systematic exposition.
-Diffeomorphisms are defined as smooth maps being also homeomorphisms.
+Diffeomorphisms are smooth maps with no singular points (of them) and being homeomorphisms.
 The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself.
-It is also topologized with the $C^{\infty}$ Whitney topology as before of course.
+It is also topologized with the $C^{\infty}$ Whitney topology as before.
 A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group.
 A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
-We present a theorem on fundamental structures of special generic maps.
-Using the notation, we also present our main result as Main Theorem 1.
-This is proven in the next section with additional results Main Theorems 2 and 3.
+We present a fundamental theorem on structures of special generic maps as Proposition 1.
+Using the notation, we present our new real algebraic maps as Main Theorem 1.
+This is proven in the next section with our additional result Main Theorems 2 and 3.
+Main Theorem 2 is a kind of remarks on Main Theorem 1.
 Main Theorem 3 is a more generalized result.
-The third section is on meanings of our main result and related topics.
+The third section is on meanings of our achievement and related topics.
 Proposition 1 ([44]).
@@
   1. (1)
-    We have an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ in such a way that $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
+    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ and $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
   2. (2)
-    We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$.
+    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$.
 By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
@@
 Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists.
-Then we have a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps on two suitably obtained mutually homeomorphic $m$-dimensional closed and connected manifolds $M$ and $M_{0}$ with the following properties.
+Then there exist a pair $(M,M_{0})$ of two mutually diffeomorphic $m$-dimensional closed and connected manifolds and a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps there and the following properties are enjoyed.
   1. (1)
-    $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+    The map $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+The map $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ being also the zero set of a real polynomial.
   2. (2)
-    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ which is also the zero set of a real polynomial.
-  3. (3)
-    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$.
+    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f_{0}(S_{0}(f_{0})))=f(S(f))$.
 This paper is regarded as an improved version of [19] and based on our new discovery and idea.
 Conflict of interest.
-This work was also supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”.
+This work has been supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”.
 The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/).
@@
 2. On Main Theorems.
-We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00. This is a main theorem of [43].
-In [19], followed by [16], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example.
+For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, let the closed interval denoted by $[a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}\leq x\leq a_{2}\}$ as is the standard notation.
+Let the open interval denoted by $(a_{1},a_{2}):={\rm Int}\ [a_{1},a_{2}]$ as is the standard case.
+Similarly, we also define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$, for example and for $a\in\mathbb{R}$, let us define $[a,\infty):=\{x\in\mathbb{R}\mid x\geq a\}$ for example.
+We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00.
+This is a main theorem of [44].
+In [20], followed by [17], the author has redefined several notions and consider specific reconstruction.
+We also abuse these notions for example.
 Proposition 2 ([44]).
@@
 We can also abuse this in the case $m=n$ and we do.
-  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
-We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
-  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
+  • Consider a small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
+We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$ and this is the first object.
+  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$.
+We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$ and this is the second object.
   • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$.
@@
 We consider the identification between their base spaces in the natural way.
-More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$.
-Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as $\partial\bar{N}\times\{1\}$ naturally and the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$.
-The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
+More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $(\partial\bar{N})\times[0,1]$.
+Furthermore, each point $x\in\partial\bar{N}$ is identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$.
+In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}$ is identified with the space $(\partial\bar{N})\times\{1\}$ naturally and it is the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$.
+The third object is a bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
 We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
-We have an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree.
-Furthermore, we have this enjoying the following properties.
+Then there exist an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree.
+Furthermore, we have the manifold and the map enjoying the following properties.
   1. (1)
-    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
+    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})=W_{f}-{\rm Int}\ N(\partial W_{f})$ in Proposition 1 are isomorphic and naturally identified.
   2. (2)
@@
 We consider their natural identification.
-This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
-Note that in Proposition 1, we cannot set ”$m\geq n$”. This is closely related to Eliashberg’s celebrated theory [6]. See also [7].
+This coincides with the identification bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ for ${\bar{f}}_{N}$.
+Note that in Proposition 1, we cannot set ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].
 Definition 1.
@@
 [20] has first introduced such a class of special generic maps.
-We first name the class in this way first in [16].
+We first name the class in this way in [17].
 Of course in our present paper, we do not need to understand this notion or [17, 20] well.
@@
 We choose the connected sum in the smooth category.
-We can have a product-organized special generic map there into ${\mathbb{R}}^{n}$ from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category.
+We can reconstruct a product-organized special generic map into ${\mathbb{R}}^{n}$ there from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category.
 This gives another simplest example.
-We add the following as another additional main result.
+We add the following new result. This is also a kind of remarks on Main Theorem 1.
 Main Theorem 2.
@@
 STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
-This is a new ingredient in our related studies. Our arguments here thanks to [20]. Especially, ”Special Case 5” and ”Definition 14”.
+This is a new ingredient in our related studies. Our arguments here thanks to [21]. Especially, ”Special Case 5” and ”Discussion 14”.
 By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$.
@@
 Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$.
-We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
-In addition, we can also get the embedding and the subset in such a way that the following are satisfied.
-  • ${\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}\geq 0$.
-  • ${\rm Int}\ {\bar{N}}_{0}$ is the intersection of the sets of all points satisfying $f_{{\partial}_{j}}>0$.
-  • At each point of ${\mathbb{R}}^{n}-{\bar{N}}_{0}$, $f_{{\partial}_{j}}<0$ holds for exactly one $1\leq j\leq l$ and $f_{{\partial}_{j}}>0$ holds for the remaining polynomials.
+We may also regard that the manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
+In addition, we can also get the embedding and the subset with the following properties.
+  • ${\bar{N}}_{0}={\bigcap}_{j=1}^{l}\{x\in{\mathbb{R}}^{n}\mid f_{{\partial}_{j}}(x)\geq 0\}$.
+  • ${\rm Int}\ {\bar{N}}_{0}={\bigcap}_{j=1}^{l}\{x\in{\mathbb{R}}^{n}\mid f_{{\partial}_{j}}(x)>0\}$.
+  • At each point $x$ of ${\mathbb{R}}^{n}-{\bar{N}}_{0}$, the relation $f_{{\partial}_{j_{x}}}(x)<0$ holds for exactly one integer $1\leq j_{x}\leq l$ and the relation $f_{{\partial}_{j}}(x)>0$ holds for the remaining polynomials.
 STEP 2 Arguments from [15, 19] to have our smooth real algebraic map $f_{0}$.
 We apply arguments from main ingredients of [15, 19].
-We consider the zero set $\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
-The last part of STEP 1 implies that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
-Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
-Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
-By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value.
-We explain about this more precisely.
-If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$.
+We consider the zero set $S_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
+The last part of STEP 1 implies the fact that $S_{0}=\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
+We apply the implicit function theorem to see that this set is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
+Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
+Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
+By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a value which is not 00.
+We discuss this more precisely.
+If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ is not 00 for some variable $x_{j_{0}}$, since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary.
+Remember that this is also obtained via an isotopy in the $C^{1}$ class.
+For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$.
 This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials.
-Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$.
-The value of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ for $x_{j_{0}}$ is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class.
-The product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$ is not zero by the property of the zero sets of the polynomials.
-We apply the implicit function theorem to see that
-$M_{0}:=\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$
-is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
-Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$.
-The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
+Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$.
+The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is not 00.
+We have checked that the set $S_{0}$ is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
+Put $M_{0}:=S_{0}$.
+Our desired map $f_{0}$ is defined as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
 STEP 3 Is $f_{0}$ special generic?
-We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $\partial{\bar{N}}_{0}\times[0,1]\subset N(\partial{\bar{N}}_{0})$.
-More precisely, we can have one and consider identification in the following way thanks to arguments related to ”Definition 14” [20].
-  • $\partial{\bar{N}}_{0}\times\{0\}$ is identified with $\partial N(\partial{\bar{N}}_{0})\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
-  • $\partial{\bar{N}}_{0}\times\{t_{0}\}\subset\partial{\bar{N}}_{0}\times(0,1]$ is identified with the zero set of the polynomial $f_{j}-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$.
-Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function: some $y_{j}$ is regarded as the variable of the target of the function.
-We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin.
+We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$.
+More precisely, we can have one and consider our identification in the following way thanks to arguments related to ”Discussion 14” in [21].
+  • The manifold $(\partial{\bar{N}}_{0})\times\{0\}$ is identified with $(\partial N(\partial{\bar{N}}_{0}))\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
+  • The manifold $(\partial{\bar{N}}_{0})\times\{t_{0}\}\subset(\partial{\bar{N}}_{0})\times(0,1]$ is identified with the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}})-a_{0}t_{0}$ for a sufficiently small number $a_{0}>0$.
+We have ${\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=at_{0}={\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ at each point $(x,y)\in(((\partial{\bar{N}}_{0})\times\{t_{0}\})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}\subset({\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ ($y=(y_{1},\cdots,y_{m-n+1})$).
+This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of a Morse function for a natural height of an ($m-n+1$)-dimensional disk and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
+Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function.
+Some $y_{j}$ is regarded as the variable of the target of the function here.
+We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin.
 This is a diffeomorphism.
-We map each point of the form $(x,0)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
-Thus we have a diffeomorphism from ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ onto ${\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ and by the restriction another diffeomorphism $({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto ${\rm Int}\ {\bar{N}}_{0}\times S^{m-n}\subset{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
+We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
+Thus we have a diffeomorphism from $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ onto $({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
+By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
 These arguments imply that $f_{0}$ is a product-organized special generic map.
@@
 This completes the proof.
-Main Theorem 3.
-Suppose a situation as in Main Theorem 1 is given.
-Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following.
-We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $t$.
-  • $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
-  • $F_{0,S^{m-n}}(0)=0$. $F_{0,S^{m-n}}(t)>0$ for $t>0$ and $F_{0,S^{m-n}}(t)<0$ for $t<0$. The value of the derivative of $F_{0,S^{m-n}}$ at 00 is not zero.
-  • ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$. $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)\geq 0$ for $y\in{\mathbb{R}}^{m-n+1}$.
-  • ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the empty set for $t<0$.
-  • For $a_{1},a_{2}\in\mathbb{R}$ such that $a_{1}<a_{2}$, define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$.
-We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)$ mapping ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively.
-${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
-  • For any real number $\epsilon>0$, we have a positive real number $\delta>0$ and for $0<t<\delta$ and each point $p$ of ${F_{S^{m-n},t}}^{-1}(0)$, $0<||p||<\epsilon$.
-Then we can have a statement of Main Theorem 1 by defining
-$M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
-for a sufficiently large $T>0$.
-As a more specific case, we can also have $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
-  1. (1)
-    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
-We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$.
-This is also a bundle which may not be smooth and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$.
-  2. (2)
-    Under the presented structure group, the bundle of (1) here which may not be a smooth bundle is in fact trivial.
-  3. (3)
-    $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps.
-    1. (a)
-      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is homeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$.
-Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4,5$.
-Note also that for $k\neq 4$ this is diffeomorphic to $D^{k}$ if its boundary is diffeomorphic to $S^{k-1}$.
-    2. (b)
-      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times D^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and $D^{m-n+1}$: here $D^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
-    3. (c)
-      Eliminate the corner.
-We review several known facts on topologies and differentiable structures of spheres and disks shortly.
+We show Main Theorem 3. This is a kind of generalizations of Main Theorem 2.
+Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly.
 See [35] and see also [13, 47, 48] for example.
@@
 If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$ or $5$.
+Main Theorem 3.
+Suppose that a situation same as one of Main Theorem 1 is given.
+Let $F_{S^{m-n}}:{\mathbb{R}}^{m-n+2}={\mathbb{R}}^{m-n+1}\times\mathbb{R}\rightarrow\mathbb{R}$ be a function defined canonically by a real polynomial enjoying the following.
+We naturally define $F_{S^{m-n},t}:{\mathbb{R}}^{m-n+1}\rightarrow\mathbb{R}$ as the restriction to ${\mathbb{R}}^{m-n+1}\times\{t\}$ for $t\in\mathbb{R}$ where ${\mathbb{R}}^{m-n+1}\times\{t\}$ and ${\mathbb{R}}^{m-n+1}$ are identified by the map mapping $(y,t)$ to $y$.
+  1. (1)
+    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
+  2. (2)
+    $F_{0,S^{m-n}}(0)=0$. $F_{0,S^{m-n}}(t)>0$ for $t>0$ and $F_{0,S^{m-n}}(t)<0$ for $t<0$. The value of the derivative of $F_{0,S^{m-n}}$ at 00 is not zero.
+  3. (3)
+    ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$. $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)\geq 0$ for $y\in{\mathbb{R}}^{m-n+1}$.
+  4. (4)
+    We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively.
+The zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the polynomial is non-singular and it does not contain the origin $0\in{\mathbb{R}}^{m-n+1}$.
+It is also homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
+Then we can have a statement of Main Theorem 1 by defining
+$M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
+for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2.
+Explicitly, we can also have the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
+  1. (5)
+    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
+We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$.
+Furthermore, the smooth bundle is a trivial smooth bundle.
+  2. (6)
+    The non-singular real algebraic manifold $M_{0}$ bounds a compact and bounded submanifold $W_{0}$ of dimension $m+1$ in ${\mathbb{R}}^{m+1}$ homeomorphic to a manifold obtained in the following steps.
+Here ${\Delta}^{k}$ is for a $k$-dimensional smooth manifold homeomorphic to $D^{k}$ ($k\geq 1$).
+This is diffeomorphic to $D^{k}$ for $k\neq 4,5$.
+This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if the boundary is diffeomorphic to $S^{k-1}$.
+    1. (a)
+      Prepare the product $\bar{N}\times{\Delta}^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$.
+    2. (b)
+      Attach the previous two manifold by a bundle isomorphism between the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}$ over $\partial\bar{N}$ and the trivial smooth bundle $(\partial\bar{N})\times{\Delta}^{m-n+1}\subset(\partial\bar{N})\times\partial{\Delta}^{m-n+2}$ over $\partial\bar{N}$ represented as the product of the identity maps on $\partial\bar{N}$ and ${\Delta}^{m-n+1}$: here ${\Delta}^{m-n+1}$ is smoothly embedded in $\partial{\Delta}^{m-n+2}$.
+    3. (c)
+      Eliminate the corner.
+Hereafter, for a real number $t$ and $x\in{\mathbb{R}}^{k}$, we use the notation respecting the structure of the canonical vector space for ${\mathbb{R}}^{k}$ such as $tx\in{\mathbb{R}^{k}}$.
+The vector $tx$ is also a point of ${\mathbb{R}}^{k}$ of course.
 A proof of Main Theorem 3.
-We can argue as STEP 1 of the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
-We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than $a$. We choose another positive number $T_{0}<a$.
-We present an argument similar to one of STEP 2 of the proof of Main Theorems 1 and 2.
-Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
-This mainly comes from the following assumption: ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
-Around a point of the form $(x,y)\in(\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 of the proof of Main Theorems 1 and 2.
-The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not zero from the assumption, and the value in the case of Main Theorems 1 and 2.
-The resulting value is not zero.
-We can argue similarly to prove that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
-We present an argument similar to one of STEP 3 of the proof of Main Theorems 1 and 2.
-For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$ and regard this as $\partial{\bar{N}}_{0}\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $\partial{\bar{N}}_{0}\times\{0\}$ are identified by the correspondence $x\mapsto(x,0)$.
+We can argue as STEP 1 in the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
+We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$.
+We choose another positive number $T_{0}<a$.
+We define $M_{0}$ as presented in the statement of Main Theorem 3.
+We also define the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ as the composition of the canonical embedding to ${\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}$ as in Main Theorems 1 and 2 and their proof.
+We present arguments similar to ones of STEP 2 in the proof of Main Theorems 1 and 2. Here we also respect the assumptions (1)–(4) here.
+First we also have the relation
+$M_{0}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
+by our definitions similarly. We show that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
+Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
+We discuss this more precisely.
+We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$.
+According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
+It is assumed in (2) that $F_{0,S^{m-n}}(t)>0$ for $t>0$.
+We can know that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
+Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 in the proof of Main Theorems 1 and 2.
+The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the value in the case of Main Theorems 1 and 2.
+The value in the case of Main Theorems 1 and 2 means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the variable $x_{j_{0}}$.
+The resulting value is not 00.
+We can argue similarly to see that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
+We present an argument similar to one in STEP 3 in the proof of Main Theorems 1 and 2.
+For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$.
+For example, we can consider this as the collar neighborhood of STEP 3 in the proof.
+We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
 We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
-First we map $\partial{\bar{N}}_{0}\times\{0\}$ to $\partial{\bar{N}}_{0}\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$ mapping $(x,0)$ to $(x,0)$.
+First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$.
+We use the natural identification mapping $(x,0)$ to $(x,0)$ here.
+According to the property (3) for example, we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
 Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
@@
   • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
-We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
+We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
 $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$
 according to the following rule.
-  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
-  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
-We have a homeomorphism onto a smooth manifold $M$.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
+  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
+Thus we have a homeomorphism from the smooth manifold $M_{0}$ onto a smooth manifold $M$. The restriction of this to the set $M_{0}-(({\partial\bar{N_{0}}})\times\{0\})$ gives a diffeomorphism.
 We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
 It is also product-organized.
-We define $W_{0}$ as
-$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
-and this can be also defined as
+We define a special generic map $f$ as this map.
+An essentially same argument also appears in Main Theorems 1 and 2 and their proof.
+We define a new set $W_{0}\subset{\mathbb{R}}^{m+1}$ as
+$\{(x,y)=(x,y_{1},\cdots y_{m-n+1})\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$
+and by our assumption, this can be also defined as
 $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid 0\leq t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T},F_{S^{m-n}}(y,t)=0\}$.
@@
 We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
-This is a diffeomorphism.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
+  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
+This rule is in fact same as the previous one.
+This gives a diffeomorphism.
 We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
-We thus have a homeomorphism from $W_{0}$.
-The image of the homeomorphism is the union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
-  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
-  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
-From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
-We can define another non-singular real algebraic manifold
-$M_{0,t}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t)=0\}$
-for $0<t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. This is diffeomorphic to $M_{0}$. This new set has properties similar to important ones of $M_{0}$.
-We go back to the trivial bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle.
+Thus we have a homeomorphism from $W_{0}$.
+The image of the homeomorphism is the disjoint union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
+From this we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
+The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$.
+In other words, it is also regarded as a ”canonical projection of $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ onto ${\rm Int}\ {\bar{N}}_{0}$”.
+We can define another set
+$M_{0,t^{\prime}}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$
+for $0<t^{\prime}\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$.
+We present several properties based on our presented arguments.
+We can also know $M_{0,t^{\prime}}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$.
+For distinct real numbers $t^{\prime}=t_{1},t_{2}$ here, the sets $M_{0,t_{1}}-((\partial\bar{N_{0}})\times\{0\})$ and $M_{0,t_{2}}-((\partial\bar{N_{0}})\times\{0\})$ are disjoint.
+These sets $M_{0,t^{\prime}}-((\partial\bar{N_{0}})\times\{0\})$ are $m$-dimensional smooth manifolds with no boundaries.
+They are seen as the graphs of smooth functions as $M_{0}-((\partial\bar{N_{0}})\times\{0\})$ and diffeomorphic to $M_{0}-((\partial\bar{N_{0}})\times\{0\})$.
+Here some variable $y_{j}$ is regarded as the variable of the target of the function.
+We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$.
+We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold.
+This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$.
+From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}})-((\partial{\bar{N}}_{0})\times\{0\}))$.
+We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle.
 This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
-In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, this smooth bundle which may not be trivial in general is also trivial.
+For this, first, remember that the projection is the restriction of the canonical projection ${\pi}_{m+1,n}$.
+Remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function.
+Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example.
+The relation $S(f_{0})=(\partial{\bar{N}}_{0})\times\{0\}$ also holds.
+In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle.
 Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
-We investigate the bundle in Main Theorem 3 (1) and (2).
-We can present a simple and natural proof.
-However, we present a complicated proof to respect general cases.
-Let the bundle in Main Theorem 3 (1) and (2) denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$.
-First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
-We apply an argument used to show Proposition 1 in a revised style.
-More precisely, Proposition 2.1 of [43] is a corresponding result.
-It also refers to [36] there.
-We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$.
-From $W_{0}$, we remove the total space of this bundle over ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$.
-We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$.
-This fiber is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$.
-This is also thanks to Ehresmann’s fibration theorem (or a relative version of this).
-The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$.
-The structure group of the bundle $M_{{\rm B},0}$ is naturally regarded as the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$.
-For each $0<t\leq 1$, it is a subbundle of the following bundle: the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly.
-For each two obtained smooth bundles for $0<t\leq 1$, fibers are regarded as a manifold homeomorphic to $\partial D^{m-n+1}$ and a manifold homeomorphic to $D^{m-n+1}$, respectively.
-The former bundle is also obtained by restricting the fiber to the boundary homeomorphic $\partial D^{m-n+1}$ of course.
-The manifold homeomorphic to $D^{m-n+1}$ is diffeomorphic to $D^{m-n+1}$ if $m-n+1\neq 4,5$ or the boundary is diffeomorphic to $\partial D^{m-n+1}=S^{m-n}$ with the constraint $m-n+1\neq 4$ of course.
-We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before.
-This element can be regarded as an element canonically extended to an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
-By these arguments, we may discuss the bundles and the structure groups in the following simple situation essentially.
-We can regard the structure group of the bundle $W_{1,t}$ as a group acting on the unit disk $D^{m-n+1}$ preserving the distance of each point and the origin $0\in D^{m-n+1}$.
-We explain about an element ${\psi}_{t}$ of the group.
-We can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}-\{0\}$ and $0<s\leq 1$.
-Of course 00 here is the origin and $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors.
-Here we may regard the relation $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$ by considering the distance $||x||$ for $x\in D^{m-n+1}-\{0\}\subset{\mathbb{R}}^{m-n+1}$.
-We consider the limit at $s=0$.
-For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$.
-Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”.
-This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
-These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1).
-We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group.
-This completes the proof of Main Theorem 3 (2).
-By our arguments, we can also check Main Theorem 3 (3) easily.
+Last we consider a small isotopy as in the proof of Main Theorems 1 and 2. This completes the proof of the statement of Main Theorem 1.
+We can also check the property (5) by investigating arguments of STEP 3 of our proof of Main Theorems 1 and 2 and ”Discussion 14” of [21] again.
+In other words, we can check that around the set $S(f_{0})$, $f_{0}$ is represented as the product map of a natural function and the identuty map on $S(f_{0})$ as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
+We can also check Main Theorem 3 (6) easily thanks to our arguments.
 A result close to this is also presented in Theorem 1 (3) of [18].
-”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there.
+There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$.
 The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there.
@@
 For this we can choose $a$ as an arbitrary positive real number.
-The bundles in (1) and (2) are regarded as ”linear bundles defined before”.
-Furthermore, in this case, we can consider a suitable situation and regard that ${\Phi}_{F_{S^{m-n}},a}$ is replaced by a diffeomorphism, that the homeomorphism mapping each sphere is regarded as a diffeomorphism, and that (3) can be discussed in the smooth category, for example.
+The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle: ${\Delta}^{k}$ is $D^{k}$ in this case.
+Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example.
 In short, we can argue all in the smooth situation in Main Theorem 3.
 This is also for Main Theorem 2.
+Main Theorem 2 gives one of most important specific cases.
 This is generalized for more general cases with $m-n\neq 4$ for example.
@@
 This is also related to studies on diffeomorphism groups of unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
+We also apply related arguments implicitly in some of our previous arguments in the present paper.
 See also Remark 4, in the next section, for example.
+We give additional explicit cases as Examples. They are also a kind of remarks.
 Example 2.
-Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a Morse function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$.
-We have a special generic map $f_{0}$ as in the case of Main Theorem 2.
-We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with $k_{j,1}$ being positive numbers and $k_{j,2}$ being positive integers.
+Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ satisfying the following.
+  • The function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is a Morse function.
+  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on a smooth closed manifold. The manifold is a non-singular real algebraic manifold of dimension $m-n$.
+  • The minimum of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is 00 and the condition ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$ holds.
+Example 1 is generalized here.
+We have a special generic map $f_{0}$ as in the case of Main Theorems 1 and 2.
+Example 3.
+We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example.
 See also Remark 3 of [15].
+Some examples of Example 2 are regarded as specific cases of the present cases.
 3. Meanings of our main result and remarks.
@@
 In [15, 16, 18], we consider cases where polynomials are given beforehand.
-In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark.
+In short, only STEP 2 in our proof of Main Theorems 1 and 2 is considered and related short exposition on STEP 3 in our proof of Main Theorems 1 and 2 is presented where the latter presentation is a kind of short remark.
 This does not affect main ingredients of [15, 16, 18].
@@
 They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
-Related to Remark 3, we add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [25, 26, 27, 28, 29, 30, 31, 32].
+Remark 3 is a remark on STEP 1 in our proof of Main Theorems 1 and 2.
+This is also independent of our present study.
+Here we also add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33].
 See also [42].
-This is also closely related to [4].
-This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory and this mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs.
-We do not explain about elementary notions on graphs precisely.
+Remark 3 is also closely related to [4].
+This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory.
+This mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs.
+We do not present elementary notions on graphs precisely.
 We except related knowledge.
 We also note that we omit several rigorous discussions in Remark 3 and that for our main ingredients of the present paper we do not need to understand this rigorously.
+However, we give comments that this is also related to Remark 1 and 2.
 Remark 3.
@@
   • We can embed $K$ by a piecewise smooth embedding $e:K\rightarrow{\mathbb{R}}^{2}$.
-  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is an smooth embedding.
+  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is a smooth embedding.
+  • At distinct vertices, the values of ${\pi}_{2,1}\circ e$ are always distinct.
   • The degree of each vertex is $1$ or $3$.
@@
 An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves.
-The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of connected components of all preimages of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$ under the natural equivalence relation.
-Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$.
-The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ containing $e(K)$ and its Poincaré-Reeb graph is isomorphic to and naturally regarded as $e(K)$.
-Moreover, we can have the situation in the following.
+The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation is considered.
+Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$.
+The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
   1. (1)
-    The restriction of ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
+    $e(K)\subset R_{K}$.
   2. (2)
+    The Poincaré-Reeb graph of $R_{K}$ is isomorphic to and naturally regarded as $e(K)$.
+  3. (3)
+    We also have the situation in the following. Of course $\overline{R_{K}}$ is the closure of $R_{K}$ in ${\mathbb{R}}^{2}$.
+    1. (a)
+      The restriction of ${\pi}_{2,1}$ or ${\pi}_{2,1}{\mid}_{\overline{R_{K}}}:\overline{R_{K}}\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
+    2. (b)
       $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
-This is also obtained by approximation of the Weierstrass type in the class $C^{2}$.
-According to arguments similar to one in ”Special Case 5” of [20], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
+A desired region $R_{K}$ can be also obtained by an approximation of the Weierstrass type in the class $C^{2}$.
+According to arguments similar to ones in ”Special Case 5” of [21], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
 The following is a remark on Main Theorem 3.
 Remark 4.
-As another method, we can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively.
+We can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively.
 In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
-We give a short remark on [17] last.
+The following is closely related to one of central problems in real algebraic geometry. This may be closely related to our future work.
 Remark 5.
-[17] has no essential errors and we have also found several more rigorous arguments.
+Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces has been difficult and important in real algebraic geometry.
+This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools.
+At present, even explicit construction and classifications for cases of real algebraic curves are difficult.
+Our pioneering studies of the construction respect unit spheres and canonical projections of them where we do not restrict attention to $1$-dimensional cases.
+Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.
+The author recognizes that our construction shows explicit cases of Viro’s methods.
+In short, Viro’s methods glue local hypersurfaces together.
+We can have non-singular real algebraic hypersurfaces by deforming and smoothing (Viro’s patchworking).
+In our paper we do not need to understand this theory well.
+However, to study more, we need to understand the theory more.
+We close our paper by a short remark on [18].
+Remark 6.
+[18] has no essential errors and we have also found that we need several more rigorous exposition.
 Some of these important arguments are discussed in the present paper.
-We will revise the version and submit it to the refereed journal, based on comments and our arguments and idea here.
+We will revise the version and submit our revised version to the refereed journal, based on the presented comments by editors and referees on [18] and our new arguments and idea in the present paper.
 Note also that the content of [18] is independent of the content of the present paper essentially.
+We have also found we can revise [18] in a self-contained way although we need to respect some of our paper.
+We will do so as we can.
 4. Acknowledgement.

### v6

@@
-Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres
+Some remark on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres
 Abstract.
-Morse functions with exactly two singular points on spheres and canonical projections of spheres are generalized to special generic maps.
-A special generic map is, roughly, a smooth map represented as the composition of a surjection onto a manifold with an immersion of codimension 00.
-Preimages of this surjections are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary.
+Morse functions with exactly two singular points on spheres and projections of naturally embedded spheres in the Euclidean spaces (the canonical projections of the unit spheres) are generalized to special generic maps.
+This is a smooth map which is the composition of a surjection onto a manifold with a smooth immersion of codimension 00.
+The preimage of the surjection is diffeomorphic to the unit sphere of a suitable dimension at each point of the interior of the manifold and a single point set at each point on the boundary.
 We construct real algebraic maps topologically seen as special generic maps whose images are smoothly embedded manifolds.
-Canonical projections of unit spheres are generalized here.
-We are interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry.
-Obtaining them explicitly is important and difficult.
-In the stories, we previously constructed some real algebraic maps topologically special generic maps.
+The canonical projections of the unit spheres are generalized here.
+We also constructed some real algebraic maps topologically special generic maps previously.
 We construct such maps in more general situations and give additional remarks.
+We are also interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry.
+The construction is important and difficult.
 Key words and phrases:
@@
 In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps.
-Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
+Such maps are generalized to special generic maps: a special generic map $c$ is a smooth map locally represented as a projection or by the form $c(x_{1},\cdots x_{m})=(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$.
 [5, 9, 44] are related pioneering studies.
 For Morse functions, see [41] for a pioneering study and [36] for systematic exposition on fundamental notions and arguments.
-For higher dimensional versions of theory of Morse functions, [52, 55] are pionnering studies.
 This paper also studies real algebraic geometry.
@@
 [39, 53] show classical and important studies and [2, 3, 21, 22] present fundamental notions and arguments and advanced studies, for example.
-Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of) spheres.
+Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of the unit) spheres.
 Celebrated theory by Nash, followed by Tognoli, says that smooth closed manifolds are non-singular real algebraic manifolds.
@@
 For example, [25, 40, 51] show explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces.
-(Unit) spheres are simplest symmetric spaces.
-We consider another generalization for projections of spheres: constructing real algebraic maps topologically special generic maps.
-The author has succeeded in such construction in [15], followed by [18] by the author himself.
+The unit spheres are simplest symmetric spaces.
+We consider another generalization for the projections of spheres: constructing real algebraic maps topologically special generic maps.
+The author has succeeded in such construction in [15], followed by [18] by the author.
 We present new related result.
@@
 A topological manifold is well-known to be homeomorphic to some CW complex.
-A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold.
+A smooth manifold is well-known to be a polyhedron canonically, and a so-called PL manifold.
 It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely.
@@
 It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
-Let $c:X\rightarrow Y$ be a differentiable map.
-$x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$.
-$c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$.
+Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$.
+$x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$ and $c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$.
 Let $S(c)$ denote the set of all singular points of $c$.
-Here, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth.
-A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$.
-This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$.
-The canonical projection of the unit sphere $S^{k-1}$ is its restriction.
-Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map.
+Here, most of differentiable maps are smooth, equivalently, of the class $C^{\infty}$.
+The canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$.
+This maps each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to $x_{1}\in{\mathbb{R}}^{k_{1}}$ with $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$.
+The canonical projection of the unit sphere $S^{k-1}$ is its restriction, rigorously.
+Here, a real algebraic manifold is the union of some connected components of the zero sets of a real polynomial map.
 Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps.
-The real affine space and the unit sphere give simplest examples.
-Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections.
-We may say Nash manifolds and maps for example for such maps.
-[46] shows related fundamental theory, arguments and some advanced theory.
+The real affine space and the unit sphere are of simplest examples.
+Our real algebraic maps are the compositions of the canonical embeddings into the real affine spaces with the canonical projections.
+Note that they may be so-called Nash manifolds and maps, as in [39].
+However, we do not use such terminologies.
 1.2. Our main result and the structure of our paper.
 Spaces of smooth maps between two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies.
-In short two smooth maps are close if their differentials of arbitrary degrees are close.
-$C^{r}$ Whitney topologies are also considered as weaker topologies for each non-negative integer $r$.
-They respect differentials of degrees at most $r$.
+Roughly, two smooth maps are close if their differentials of arbitrary degrees are close.
+$C^{r}$ Whitney topologies are weaker topologies, respecting differentials of degrees at most $r$.
 The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$.
-Such topologies are important in singularity theory and differential topology.
-We omit rigorous exposition.
-See [11] for systematic exposition.
-Diffeomorphisms are smooth maps with no singular points (of them) and being homeomorphisms.
-The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself.
-It is also topologized with the $C^{\infty}$ Whitney topology as before.
-A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group.
-A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
-We present a fundamental theorem on structures of special generic maps as Proposition 1.
-Using the notation, we present our new real algebraic maps as Main Theorem 1.
-This is proven in the next section with our additional result Main Theorems 2 and 3.
-Main Theorem 2 is a kind of remarks on Main Theorem 1.
-Main Theorem 3 is a more generalized result.
-The third section is on meanings of our achievement and related topics.
+These topologies are important in singularity theory and differential topology.
+Rigorous exposition on singularity theory of differential maps including these topologies is left to [11].
+Diffeomorphisms are smooth homeomorphisms with no singular points (of them).
+The diffeomorphism group of a smooth manifold is the group of all diffeomorphisms from the manifold onto itself.
+A smooth bundle is a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group.
+A linear bundle is a smooth bundle whose fiber is the Euclidean space, the unit disk, or the unit sphere and whose structure group consists of linear transformations, defined naturally.
+Proposition 1 explains a fundamental theorem on structures of special generic maps.
+Using the notation, we present our new real algebraic map which is special generic topologically in Main Theorem 1, proven in the next section with our additional result Main Theorems 2 and 3.
+Main Theorem 2 is for additional comments on Main Theorem 1.
+Main Theorem 3 generalizes them.
+The third section presents additional comments on our new result.
 Proposition 1 ([44]).
 Let $m>n\geq 1$ be integers.
-For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
+For a special generic map $f:M\rightarrow N$ from an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
+Note that the case $N:={\mathbb{R}}^{n}$ is mainly discussed in [44] and our paper, and that as is also presented shortly in [44], $N$ can be general.
   1. (1)
-    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ and $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
+    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$, mapping $S(f)$ onto $\partial W_{f}$ by a diffeomorphism, and a smooth immersion $\bar{f}$.
   2. (2)
-    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$.
-By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
+    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$.
+By (the restriction of) $q_{f}$, we also have a smooth bundle over the closure of $W_{f}-N(\partial W_{f})$ in $W_{f}$ whose fiber is the unit sphere $S^{m-n}$.
 Main Theorem 1.
@@
 Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists.
-Then there exist a pair $(M,M_{0})$ of two mutually diffeomorphic $m$-dimensional closed and connected manifolds and a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps there and the following properties are enjoyed.
+Then we have an $m$-dimensional non-singular real algebraic manifold $M$ being also compact, connected, and the zero set of a real polynomial function and the following maps.
   1. (1)
-    The map $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
-The map $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ being also the zero set of a real polynomial.
+    A special generic map $f:M\rightarrow N:={\mathbb{R}}^{n}$ in Proposition 1 with $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+This map $f$ is not real algebraic in general.
   2. (2)
-    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f_{0}(S_{0}(f_{0})))=f(S(f))$.
-This paper is regarded as an improved version of [19] and based on our new discovery and idea.
+    A real algebraic map $f_{0}:M\rightarrow{\mathbb{R}}^{n}$ and a homeomorphism $\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f_{0}=\phi\circ f$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$: remember that $S(c)$ denotes the singular set of a smooth map $c$.
+The simplest case $\bar{N}=D^{n}$ accounts for the canonical projection $f=f_{0}:S^{m}\rightarrow{\mathbb{R}}^{n}$ of the unit sphere $S^{m}$ with $\phi$ being the identity map.
 Conflict of interest.
-This work has been supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”.
+This work was also supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”.
 The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/).
@@
 Data availability.
-Data essentially supporting our present study are all here.
+Data essentially supporting our present study are all here. This paper is regarded as an improved version of one of a preprints of the author [19].
 2. On Main Theorems.
@@
 Similarly, we also define $(a_{1},a_{2}]:=\{x\in\mathbb{R}\mid a_{1}<x\leq a_{2}\}$, for example and for $a\in\mathbb{R}$, let us define $[a,\infty):=\{x\in\mathbb{R}\mid x\geq a\}$ for example.
-We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00.
-This is a main theorem of [44].
+As an important ingredient in special generic maps, we explain the Morse function for the natural height of the disk $D^{k}\subset{\mathbb{R}}^{k}$.
+We map $x\in D^{k}$ to $(\sqrt{1-{||x||}^{2}},x)\in\mathbb{R}\times D^{k}\subset{\mathbb{R}}^{k+1}$ by a diffeomorphism, compose the projection to the second component and we have the Morse function for the natural height of the disk $D^{k}\subset{\mathbb{R}}^{k}$.
+The projection of $D^{k}$ embedded in ${\mathbb{R}}^{k+1}$ here is the restriction of ${\pi}_{k+1,1}{\mid}_{S^{k}}$ to $S^{k}\bigcap\{x:=(x_{1},\cdots x_{k+1})\mid x_{1}\geq 0\}$.
+Locally, a special generic map from an $m$-dimensional manifold with no boundary into an $n$-dimensional manifold with no boundary with $m\geq n$ is a smooth map such that around each singular point of it, the map is represented as the product map of such a function and the identity map on an ($n-1$)-dimensional manifold (where suitable local coordinates are chosen).
+Related to Proposition 1, we also review a theorem reconstructing special generic maps from given smooth immersions of codimension 00, conversely.
+Note again that the case $N:={\mathbb{R}}^{n}$ is discussed in the original study mainly, and that we can consider the general case.
 In [20], followed by [17], the author has redefined several notions and consider specific reconstruction.
@@
 Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given.
-Hereafter we can abuse the notation in Proposition 1 and do.
+Hereafter, we can abuse the notation in Proposition 1 and do.
 We can also abuse this in the case $m=n$ and we do.
-  • Consider a small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
-We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$ and this is the first object.
-  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$.
-We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$ and this is the second object.
-  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$.
+  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over the closure $\overline{\bar{N}-N(\partial\bar{N})}$ of $\bar{N}-N(\partial\bar{N})$ in $\bar{N}$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$.
+We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
+  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
+  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\overline{\bar{N}-N(\partial\bar{N})})$ of the closure $\overline{\bar{N}-N(\partial\bar{N})}$.
 We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$.
@@
 More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $(\partial\bar{N})\times[0,1]$.
-Furthermore, each point $x\in\partial\bar{N}$ is identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$.
-In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}$ is identified with the space $(\partial\bar{N})\times\{1\}$ naturally and it is the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$.
-The third object is a bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$.
-We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
+Furthermore, each point $x\in\partial\bar{N}$ is naturally identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$, of course.
+In addition, in considering the base space of ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial N(\partial\bar{N})-\partial\bar{N}=(\partial\bar{N})\times\{1\}$ as $\partial\bar{N}$, $x\in\partial\bar{N}$ and $(x,0)\in(\partial\bar{N})\times\{0\}$ are also seen as $(x,1)\in(\partial\bar{N})\times\{1\}$.
+    In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}=(\partial\bar{N})\times\{1\}$ is also regarded as the subspace of $\overline{\bar{N}-N(\partial\bar{N})}\subset\bar{N}$ and these two spaces are naturally identified.
+The third object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ over $\partial(\overline{\bar{N}-N(\partial\bar{N})})\subset\overline{\bar{N}-N(\partial\bar{N})}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$ over $\partial\bar{N}=\partial N(\partial\bar{N})-\partial\bar{N}$ where the induced map between the base spaces is the natural identification for the spaces of course: note also that we identify several spaces according to the identification before.
+We call this bundle isomorphism an identification bundle isomorphism for ${\bar{f}}_{N}$.
 Then there exist an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree.
-Furthermore, we have the manifold and the map enjoying the following properties.
+We can also have the manifold and the map enjoying the following properties.
   1. (1)
-    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})=W_{f}-{\rm Int}\ N(\partial W_{f})$ in Proposition 1 are isomorphic and naturally identified.
+    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ in Proposition 1 are isomorphic and naturally identified via a diffeomorphism ${\Phi}_{M,{\rm I}}:{\bar{M}}_{\rm I}\rightarrow M-{q_{f}}^{-1}(N(\partial\bar{N}))$ and the identity map on $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ where the closure is taken in $\bar{N}=W_{f}$ as before.
   2. (2)
-    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified.
+    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified via a diffeomorphism ${\Phi}_{M,{\rm B}}:{\bar{M}}_{\rm B}\rightarrow{q_{f}}^{-1}(N(\partial\bar{N}))$ and the identity map on $N(\partial\bar{N})=N(\partial W_{f})$.
   3. (3)
-    We also consider the restriction of the smooth bundle in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here.
-We consider their natural identification.
-This coincides with the identification bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ for ${\bar{f}}_{N}$.
-Note that in Proposition 1, we cannot set ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].
+    We consider the bundle over $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ and the linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1.
+Their restrictions to the boundaries are identified by the identification induced canonically from the manifold $M$ and the map $f:M\rightarrow N$.
+    We also consider the restriction of the smooth bundle ${\bar{M}}_{\rm I}$ in (1) to the boundary and the subbundle of the linear bundle ${\bar{M}}_{\rm B}$ whose fiber is $\partial D^{m-n+1}$ in (2).
+We consider their identification or the bundle isomorphism induced from (1, 2) and the short arguments on $M$ and $f:M\rightarrow N$ just before: the identification coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
+Remark 1.
+This remarks reconstruction of a special generic map $f:M\rightarrow N$ in Proposition 2.
+  1. (1)
+    By considering the identification bundle isomorphism for ${\bar{f}}_{N}$ and gluing (the spaces of the domains of) ${\Phi}_{M,{\rm I}}$ and ${\Phi}_{M,{\rm B}}$ by this, we have a diffeomorphism ${\Phi}_{M,{\rm I},{\rm B}}$ onto our new manifold $M$ such that ${\Phi}_{M,{\rm I}}$ and ${\Phi}_{M,{\rm B}}$ are regarded as its restrictions.
+  2. (2)
+    We can regard that the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ naturally reconstructs a smooth surjection $q_{f}:M\rightarrow W_{f}=\bar{N}$.
+After that, we compose the given smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ to have our desired special generic map $f:M\rightarrow N$.
+Note that in Proposition 1, we cannot consider the condition ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].
 Definition 1.
-In Proposition 2, if the interior unit sphere bundle and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is given as the product map of the identity map (identification) between the base spaces and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
+In Proposition 2, if the interior unit sphere bundle for ${\bar{f}}_{N}$ and the boundary linear bundle for ${\bar{f}}_{N}$ are trivial and the identification bundle isomorphism for ${\bar{f}}_{N}$ is the product map of the identity map on the base space (or the identification between the base spaces) and the identity map on the fiber $S^{m-1}=\partial D^{m-n+1}$, then it is called a product-organized special generic map.
 [20] has first introduced such a class of special generic maps.
 We first name the class in this way in [17].
-Of course in our present paper, we do not need to understand this notion or [17, 20] well.
-We do not assume their arguments or results.
-However, we note that canonical projections of unit spheres are of such a class.
+Of course in our present paper, we do not need to understand precise arguments or results from [17, 20] well.
+We note the canonical projection of the unit sphere is of such a class.
 We consider a manifold represented as a connected sum $\sharp(S^{k_{j}}\times S^{m-k_{j}})$ of finitely many manifolds $S^{k_{j}}\times S^{m-k_{j}}$ where $m\geq n\geq 2$ and $11\leq k_{j}\leq n-$ 1 are integers.
-We choose the connected sum in the smooth category.
+We take the connected sum in the smooth category.
 We can reconstruct a product-organized special generic map into ${\mathbb{R}}^{n}$ there from a natural embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ whose image is diffeomorphic to the boundary connected sum $\natural(D^{k_{j}}\times S^{n-k_{j}})$ of finitely many manifolds $D^{k_{j}}\times S^{n-k_{j}}$ taken in the smooth category.
 This gives another simplest example.
-We add the following new result. This is also a kind of remarks on Main Theorem 1.
+We have the following by the definition and the structures of the maps immediately. In the following, we abuse the notation of Propositions 1 and 2 or its natural variant.
+Proposition 3.
+Let $f_{1}:M\rightarrow N$ be a product-organized special generic map and for the induced smooth immersion $\bar{f}:=\bar{f_{1}}:W_{f_{1}}\rightarrow N$ and another smooth immersion ${{\bar{f}}_{N}}^{\prime}:W_{f_{1}}\rightarrow N$, let there exist a homeomorphism ${\phi}_{N_{1,2}}:N\rightarrow N$ satisfying the relation ${{\bar{f}}_{N}}^{\prime}={\phi}_{N_{1,2}}\circ\bar{f_{1}}$.
+In this situation, there exists a product-organized special generic map $f_{2}:M\rightarrow N$ satisfying the relation $f_{2}={\phi}_{N_{1,2}}\circ f_{1}$ and $\bar{f_{2}}={{\bar{f}}_{N}}^{\prime}$.
+In short, Remark 1 (2) gives Proposition 3.
+We explain some, more precisely.
+First, from $f_{1}$, we have two kinds of bundles as in Proposition 1.
+From these two bundles and the identification between the boundaries, we obtain the three objects for reconstruction of a special generic map in Proposition 2 canonically.
+We can consider a smooth surjection as in Remark 1 (2), and after that, we compose the immersion ${{\bar{f}}_{N}}^{\prime}$ to have our desired special generic map $f_{2}$.
+The following new result is a kind of our remark on Main Theorem 1.
 Main Theorem 2.
@@
 STEPs 1–3 are main ingredients.
-STEP 1 Obtaining another nice compact manifold isotopic to the original embedding.
+STEP 1 Obtaining another nice compact manifold which is isotopic to the original embedding and whose boundary consists of connected components each of which is the zero set of some real polynomial function and non-singular.
 This is a new ingredient in our related studies. Our arguments here thanks to [21]. Especially, ”Special Case 5” and ”Discussion 14”.
-By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+By considering a suitable small isotopy of the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another nice smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ in the following.
+Before explaining the new embedding, we give some small remark.
 We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [23].
 For this approximation, see also [4, 8].
-This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$.
+This can be also chosen as an isotopy fixing the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$.
 Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$.
 Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$.
-We may also regard that the manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
+The manifold ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$.
 In addition, we can also get the embedding and the subset with the following properties.
@@
 For the variable $x_{j_{0}}$, the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$.
-This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials.
+This is thanks to elementary formula in calculus and the property of the zero sets of our polynomial functions.
 Remember that $f_{{\partial}_{j}}(x)$ is not 00 for any $j$ except $j=i_{0}$.
-The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ is not 00.
+The value of the partial derivative of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ (for $x_{j_{0}}$) is not 00.
 We have checked that the set $S_{0}$ is a smooth, closed and connected manifold and also a non-singular real algebraic manifold of dimension $m$.
@@
 We can have a small collar neighborhood $N(\partial{\bar{N}}_{0})$ of the boundary $\partial{\bar{N}}_{0}$, regarded as $(\partial{\bar{N}}_{0})\times[0,1]\subset N(\partial{\bar{N}}_{0})$.
-More precisely, we can have one and consider our identification in the following way thanks to arguments related to ”Discussion 14” in [21].
+More precisely, we can have one and consider our identification in the following way thanks to arguments related to [21, Discussion 14].
   • The manifold $(\partial{\bar{N}}_{0})\times\{0\}$ is identified with $(\partial N(\partial{\bar{N}}_{0}))\bigcap\partial{\bar{N}}_{0}=\partial{\bar{N}}_{0}$.
@@
 We have ${\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=at_{0}={\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ at each point $(x,y)\in(((\partial{\bar{N}}_{0})\times\{t_{0}\})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}\subset({\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ ($y=(y_{1},\cdots,y_{m-n+1})$).
-This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of a Morse function for a natural height of an ($m-n+1$)-dimensional disk and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
-Around a point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ in $M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function.
+This means that around $N(\partial{\bar{N}}_{0})$, $f_{0}$ is a special generic map and represented as the product map of the Morse function for the natural height of the ($m-n+1$)-dimensional unit disk $D^{m-n+1}$ and the identity map on $\partial{\bar{N}}_{0}$ (for suitable local coordinates).
+Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, remember that locally $M_{0}$ is represented as the graph of a canonically obtained smooth function.
 Some $y_{j}$ is regarded as the variable of the target of the function here.
-We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$ according to the following rule in the case $y$ is not the origin.
-This is a diffeomorphism.
+We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\})$ to $(x,\frac{1}{\sqrt{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}}y)$.
+This gives a diffeomorphism on $({\rm Int}\ {\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\})$.
 We map each point of the form $(x,0)\in({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
@@
 By the restriction, we also have another diffeomorphism $(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ onto $({\rm Int}\ {\bar{N}}_{0})\times S^{m-n}\subset({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1}$.
-These arguments imply that $f_{0}$ is a product-organized special generic map.
+The structures of the manifolds and maps and these arguments imply that $f_{0}$ is a product-organized special generic map.
 We argue STEPs 1–3 again in a generalized way in the proof of Main Theorem 3.
-By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}^{n}}$ enjoying all desired properties in Main Theorems 1 and 2.
+We check the existence of desired maps for Main Theorems 1 and 2.
+If we do not need to deform the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ in STEP 1 and we can regard ${\bar{f}}_{{\mathbb{R}}^{n}}={\bar{f}}_{{\mathbb{R}}^{n},0}$, then we can have desired maps by putting the special generic map $f:=f_{0}$ and the homeomorphism $\phi$ as the identity map.
+In general, we need to deform the embedding in STEP 1 and in the general case, our desired map $f$ is reconstructed by Propositions 2 and 3 and Definition 1 similarly from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$.
+For the original embedding, each connected component of the boundary of the image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$ may not be real algebraic (as a result $f$ may not be real algebraic).
+We can also know the existence of the homeomorphism $\phi$ for Main Theorems 1 and 2 by the structures of the manifolds and maps and Proposition 3, easily.
 This completes the proof.
-We show Main Theorem 3. This is a kind of generalizations of Main Theorem 2.
+We show a kind of generalizations for Main Theorems 1 and 2.
 Before this, we review several known facts on topologies and differentiable structures of spheres and disks shortly.
-See [35] and see also [13, 47, 48] for example.
+See [35] and see also [13, 47, 48], for example.
 A smooth manifold which is homeomorphic to $S^{k}$ ($k\geq 1$) is diffeomorphic to $S^{k}$ for $k=1,2,3,5,6$ for example.
@@
   1. (1)
-    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
+    $F_{S^{m-n}}(y,t)=F_{0,S^{m-n}}(t)-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some real polynomials $F_{0,S^{m-n}}$ and $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ and $(y,t)\in{\mathbb{R}}^{m-n+1}\times\mathbb{R}$.
   2. (2)
@@
   4. (4)
-    We have a diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a diffeomorphism where $a$ and ${\Sigma}^{m-n}$ are a suitably chosen positive real number, and a suitably chosen smooth manifold homeomorphic to $S^{m-n}$, respectively.
-The zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the polynomial is non-singular and it does not contain the origin $0\in{\mathbb{R}}^{m-n+1}$.
-It is also homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
-Then we can have a statement of Main Theorem 1 by defining
+    A positive number $a>0$ and a smooth manifold ${\Sigma}^{m-n}$ with a homeomorphism ${\phi}_{\Sigma}:{\Sigma}^{m-n}\rightarrow S^{m-n}$ are given.
+There exists a homeomorphism ${\Phi}_{F_{{S}^{m-n}},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a homeomorphism.
+Furthermore, the zero set ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ of the real polynomial function is non-singular for $t\in(0,a]$.
+In this situation, by defining
 $M_{0}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
-for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2.
+for a sufficiently large $T>0$ and real polynomials $f_{{\partial}_{j}}$ in our proof of Main Theorems 1 and 2, we have an $m$-dimensional smooth manifold homeomorphic to $M_{0}$ and the following maps, similar to those in Main Theorem 1, as follows.
+  1. (5)
+    A special generic map $f:M\rightarrow N:={\mathbb{R}}^{n}$ in Proposition 1 with $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$.
+  2. (6)
+    A real algebraic map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f_{0}\circ\Phi=\phi\circ f$ for some pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms with the relations $\Phi(S(f))=S(f_{0})$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$.
 Explicitly, we can also have the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ with the following properties.
   1. (7)
-    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example.
-We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and the boundary of whose fiber is homeomorphic to $S^{m-n}$.
+    We can abuse the notation in Proposition 1 replacing ”$f$” by ”$f_{0}$” for example.
+We have a smooth bundle whose projection is the composition of the restriction of $q_{f_{0}}$ to the preimage of a suitable small collar neighborhood $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, and the boundary of whose fiber is homeomorphic to $S^{m-n}$.
 Furthermore, the smooth bundle is a trivial smooth bundle.
@@
 This is diffeomorphic to $D^{k}$ for $k\neq 4,5$.
-This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if the boundary is diffeomorphic to $S^{k-1}$.
+This is also diffeomorphic to $D^{k}$ for $k\neq 4$ if its boundary is diffeomorphic to $S^{k-1}$.
     1. (a)
@@
 A proof of Main Theorem 3.
-We can argue as STEP 1 in the proof of Main Theorems 1 and 2. We can abuse the notation there and we do.
+We can argue as ”STEP 1 in the proof of Main Theorems 1 and 2”. We can abuse the notation there and we do.
 We choose a sufficiently large number $T>0$ such that the maximal value of $\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$ is smaller than the given positive number $a>0$.
-We choose another positive number $T_{0}<a$.
 We define $M_{0}$ as presented in the statement of Main Theorem 3.
 We also define the map $f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n}$ as the composition of the canonical embedding to ${\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}$ as in Main Theorems 1 and 2 and their proof.
-We present arguments similar to ones of STEP 2 in the proof of Main Theorems 1 and 2. Here we also respect the assumptions (1)–(4) here.
+We present arguments similar to ones of ”STEP 2 in the proof of Main Theorems 1 and 2”. Here we also respect the assumptions (1)–(4) here.
 First we also have the relation
 $M_{0}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}\mid F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=0\}$
-by our definitions similarly. We show that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
+by our definitions and the assumptions (1)–(3) similarly. The corresponding argument ”in the proof of Main Theorems 1 and 2” is naturally generalized.
+We check that $M_{0}$ is also a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
 Around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function.
@@
 We consider the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ for some variable $y_{j}$.
-According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular, which does not contain the origin and which is homeomorphic to $S^{m-n}$ for $t\in(0,a]$.
-It is assumed in (2) that $F_{0,S^{m-n}}(t)>0$ for $t>0$.
-We can know that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
-Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in STEP 2 in the proof of Main Theorems 1 and 2.
-The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the value in the case of Main Theorems 1 and 2.
-The value in the case of Main Theorems 1 and 2 means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the variable $x_{j_{0}}$.
+According to the property (4), ${F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ is the zero set of the real polynomial which is non-singular for $t\in(0,a]$.
+We have checked that the partial derivative of the polynomial $F_{S^{m-n}}(y,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})=F_{0,S^{m-n}}(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ is not 00 for some variable $y_{j}$.
+Around a point of the form $(x,y)\in((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, we have a partial derivative as in ”STEP 2 in the proof of Main Theorems 1 and 2”.
+The value is the product of the derivative of $F_{0,S^{m-n}}$ at 00, which is not 00 from the assumption (2), $\frac{1}{T}$, and the ”value in the case of Main Theorems 1 and 2”.
+The ”value in the case of Main Theorems 1 and 2” means the value of the partial derivative of ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))$ for the ”variable $x_{j_{0}}$ there”.
 The resulting value is not 00.
-We can argue similarly to see that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
-We present an argument similar to one in STEP 3 in the proof of Main Theorems 1 and 2.
+We have checked similarly that $M_{0}$ is a smooth closed and connected manifold and a non-singular real algebraic manifold of dimension $m$.
+We present an argument similar to one in ”STEP 3 in the proof of Main Theorems 1 and 2”.
 For the boundary of the image ${\bar{N}}_{0}$, we consider a small collar neighborhood $N(\partial{\bar{N}}_{0})$.
-For example, we can consider this as the collar neighborhood of STEP 3 in the proof.
+For example, we can consider this as the collar neighborhood of ”STEP 3 in the proof of Main Theorems 1 and 2”.
 We regard this as $(\partial{\bar{N}}_{0})\times[0,1]$ where $\partial{\bar{N}}_{0}$ and $(\partial{\bar{N}}_{0})\times\{0\}$ are identified by the correspondence identifying $x$ with $(x,0)$.
-We construct a homeomorphism from the manifold $M_{0}$ onto another smooth manifold $M$.
+We construct a homeomorphism ${\Phi}^{\prime}$ from the manifold $M_{0}$ onto another new smooth manifold $M$.
 First we map $(\partial{\bar{N}}_{0})\times\{0\}$ to $(\partial{\bar{N}}_{0})\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}$.
 We use the natural identification mapping $(x,0)$ to $(x,0)$ here.
-According to the property (3) for example, we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
-Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a,1}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component and ${\rm Pr}_{{\Sigma}^{m-n},a,2}:{\Sigma}^{m-n}\times(0,a]\rightarrow(0,a]$ the projection to the second component.
+According to the assumption (3), we have $((\partial{\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}=(\partial{\bar{N}}_{0})\times\{0\}$.
+Hereafter, let ${\rm Pr}_{{\Sigma}^{m-n},a}:{\Sigma}^{m-n}\times(0,a]\rightarrow{\Sigma}^{m-n}$ denote the projection to the first component.
 We can consider a smooth function $g:(0,\infty)\rightarrow(0,1]$ enjoying the following.
@@
   • $g(t)=1$ for $t\geq 1-\epsilon$ with a sufficiently small $\epsilon>0$.
+Remember that our original homeomorphism ${\Phi}_{F_{S^{m-n}},a}$ is presented in the assumption (4): this maps ${\Sigma}^{m-n}\times(0,a]$ onto ${\bigcup}_{t\in(0,a]}{F_{S^{m-n},t}}^{-1}(0)\subset{\mathbb{R}}^{m-n+1}$ mapping each homotopy sphere ${\Sigma}^{m-n}\times\{t\}$ onto ${F_{{S}^{m-n},t}}^{-1}(0)$ by a homeomorphism.
 We can map each point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$ to
-$(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),T_{0}))$
+$(x,g(t){\phi}_{\Sigma}\circ{\rm Pr}_{{\Sigma}^{m-n},a}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y))$
 according to the following rule.
-  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the value of the projection to the 2nd component of $x$ here.
   • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
-Thus we have a homeomorphism from the smooth manifold $M_{0}$ onto a smooth manifold $M$. The restriction of this to the set $M_{0}-(({\partial\bar{N_{0}}})\times\{0\})$ gives a diffeomorphism.
-We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
+Thus we have a continuous bijection ${\Phi}^{\prime}$ between compact and connected manifolds, which is also a homeomorphism ${\Phi}^{\prime}$ from the smooth manifold $M_{0}$ onto another smooth manifold $M$.
+We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
 The resulting map is also special generic by the construction.
 It is also product-organized.
-We define a special generic map $f$ as this map.
-An essentially same argument also appears in Main Theorems 1 and 2 and their proof.
+We define a special generic map $f^{\prime}:M\rightarrow{\mathbb{R}}^{n}$ as this map.
+Note that this is not real algebraic (in general).
+An essentially same argument also appears in Main Theorems 1 and 2 and their proof. We further present another argument, an argument similar to which also appears there.
 We define a new set $W_{0}\subset{\mathbb{R}}^{m+1}$ as
@@
 We use the function $g$ before.
-We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
-  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
+We can map each point of the form $(x,y)\in({\rm Int}\ ({\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\}))\bigcap W_{0}$ to $(x,g(t){\phi}_{\Sigma}\circ{\rm Pr}_{{\Sigma}^{m-n},a}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y))$ according to the following rule.
+  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the value of the projection to the 2nd component of $x$ here.
   • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
 This rule is in fact same as the previous one.
-This gives a diffeomorphism.
+This gives a continuous bijection.
 We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$.
-Thus we have a homeomorphism from $W_{0}$.
-The image of the homeomorphism is the disjoint union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
-From this we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
+Thus we have a continuous bijection from $W_{0}$ to another compact and connected manifold, extending the homeomorphism ${\Phi}^{\prime}$ to a new homeomorphism.
+The image of the homeomorphism is the disjoint union of the image of the homeomorphism just before and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
+For here, remember again the condition (3), assuming ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$.
+From this, we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
 The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$.
@@
 Here some variable $y_{j}$ is regarded as the variable of the target of the function.
-We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$.
 We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold.
 This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$.
-From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}})-((\partial{\bar{N}}_{0})\times\{0\}))$.
+We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$.
+From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\sqcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}}-((\partial{\bar{N}}_{0})\times\{0\})))$.
 We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle.
 This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this).
-For this, first, remember that the projection is the restriction of the canonical projection ${\pi}_{m+1,n}$.
-Remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function.
+For this, remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function.
 Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example.
-The relation $S(f_{0})=(\partial{\bar{N}}_{0})\times\{0\}$ also holds.
-In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle.
-Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
-Last we consider a small isotopy as in the proof of Main Theorems 1 and 2. This completes the proof of the statement of Main Theorem 1.
-We can also check the property (5) by investigating arguments of STEP 3 of our proof of Main Theorems 1 and 2 and ”Discussion 14” of [21] again.
-In other words, we can check that around the set $S(f_{0})$, $f_{0}$ is represented as the product map of a natural function and the identuty map on $S(f_{0})$ as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
-We can also check Main Theorem 3 (6) easily thanks to our arguments.
-A result close to this is also presented in Theorem 1 (3) of [18].
+In the case $F_{S^{m-n}}(y,t)=t-||y||^{2}$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle.
+Our arguments here generalize ones from ”STEP 3 in the proof of Main Theorems 1 and 2”.
+These arguments also imply the relation $f_{0}\circ{\Phi}^{\prime}=f^{\prime}$.
+We can also discuss for the existence of the homeomorphisms of (6) similarly.
+More precisely, to obtain a product-organized special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}$ and a homeomorphism $\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f^{\prime}=\phi\circ f$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$, we apply Proposition 3 with Remark 1 (2): remember also that the embeddings ${\bar{f}}_{{\mathbb{R}}^{n}}$ and ${\bar{f}}_{{\mathbb{R}}^{n},0}$ are mutually isotopic smooth embeddings.
+In addition, the homeomorohism ${\Phi}^{\prime}$ is our desired homeomorphism $\Phi$.
+This completes the proof of the statement similar to that of Main Theorem 1 with the properties (5, 6).
+We can also check the property (7) easily by our construction.
+We can check that around the set $S(f_{0})$, $f_{0}$ is represented as a natural product map as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
+We can also check Main Theorem 3 (8) easily by our construction.
+A result close to this is also presented in the previous version (2nd version) of [18] ([18, Theorem 1 (3)]).
 There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$.
@@
 Only exposition on ”homotopy” is presented there.
+Of course we do not assume arguments and result of [18].
 This completes the proof. ∎
 Example 1.
-Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$.
+The case $F_{S^{m-n}}(y,t)=t-||y||^{2}$ ($y\in{\mathbb{R}}^{m-n+1}$) gives one of simplest cases.
 For this we can choose $a$ as an arbitrary positive real number.
-The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle: ${\Delta}^{k}$ is $D^{k}$ in this case.
+The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle with ${\Delta}^{k}=D^{k}$ in this case.
 Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example.
-In short, we can argue all in the smooth situation in Main Theorem 3.
-This is also for Main Theorem 2.
-Main Theorem 2 gives one of most important specific cases.
+In short, we can argue all in the smooth situation in Main Theorem 3 except arguments related to the deformation of the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ of STEP 1.
+This is also for Main Theorems 1 and 2.
+Main Theorems 1 and 2 give one of most important specific and simplest cases.
 This is generalized for more general cases with $m-n\neq 4$ for example.
@@
 This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres.
-This is also related to studies on diffeomorphism groups of unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
-We also apply related arguments implicitly in some of our previous arguments in the present paper.
-See also Remark 4, in the next section, for example.
-We give additional explicit cases as Examples. They are also a kind of remarks.
+This is also related to studies on diffeomorphism groups of the unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
+We also apply related arguments implicitly in some of our previous arguments in the present paper. See also Remark 5, in the next section, for example.
+We give additional explicit cases as Examples. They are also a kind of our new remark.
 Example 2.
 Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ satisfying the following.
-  • The function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is a Morse function.
-  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on a smooth closed manifold. The manifold is a non-singular real algebraic manifold of dimension $m-n$.
+  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on an ($m-n+1$)-dimensional smooth manifold with no boundary.
   • The minimum of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is 00 and the condition ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$ holds.
+  • The preimage $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(t)$ is a non-singular real algebraic manifold of dimension $m-n$ for $t>0$ and diffeomorphic to $S^{m-n}$.
 Example 1 is generalized here.
@@
 We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example.
-See also Remark 3 of [15].
+See also [15, Remark 3].
 Some examples of Example 2 are regarded as specific cases of the present cases.
-3. Meanings of our main result and remarks.
-The first remark compares our main result to previous results by the author.
+3. Additional comments on our new result.
+Remark 2 compares our main result to previous results by the author.
 Remark 2.
-In [15, 16, 18], we consider cases where polynomials are given beforehand.
-In short, only STEP 2 in our proof of Main Theorems 1 and 2 is considered and related short exposition on STEP 3 in our proof of Main Theorems 1 and 2 is presented where the latter presentation is a kind of short remark.
+In [15, 16, 18], we consider cases where in Main Theorems, we do not need to change the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ and can regard ${\bar{f}}_{{\mathbb{R}}^{n}}={\bar{f}}_{{\mathbb{R}}^{n},0}$ with the polynomials $f_{{\partial}_{j}}$ being given beforehand.
+In short, only ”STEP 2 in our proof of Main Theorems 1 and 2” is considered and related short exposition on ”STEP 3 in our proof of Main Theorems 1 and 2” is presented there.
+The latter presentation is a kind of short remark.
 This does not affect main ingredients of [15, 16, 18].
-STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first.
-These studies also consider explicit cases.
-For example, polynomials are of degree $2$ and for spheres.
-These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
+”Our STEP 3” explicitly shows that $f_{0}$ is ”topologically special generic” first.
+These studies [15, 16, 18] also consider very explicit cases.
+For example, the polynomials are of degree $2$ and for spheres.
+These studies are also regarded as studies finding nice classes of real algebraic maps including the canonical projections of the unit spheres as simplest cases.
 The following is on another problem.
 Remark 3.
-As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing canonical projections. Some are closely related to Remark 1.
+As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing the canonical projections to the real algebraic maps in Main Theorems.
+Some are closely related to Remark 2, just before.
 More precisely, such studies are on Reeb graphs of these functions.
-The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages.
+The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of all connected components of all preimages of single points.
 The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
@@
 In 2006, [45] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph.
-[24, 34] show explicit answers.
+[24, 34] also show explicit answers.
 [14] is motivated by such studies.
-There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
-[15] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first.
-This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere.
-For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices.
+There the author has succeeded in construction of nice smooth functions with prescribed preimages, first.
+[15] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations, first.
+This is also regarded as a study finding nice classes of real algebraic functions including the canonical projection of the unit sphere into $\mathbb{R}$.
+For the Reeb graph of the canonical projection of the unit sphere of dimension at least $2$ into $\mathbb{R}$, its Reeb graph consists of exactly one edge and exactly two vertices.
 For this, [4, 49, 50] are also closely related, help the author to do such studies, and motivate us to study further.
-They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
-Remark 3 is a remark on STEP 1 in our proof of Main Theorems 1 and 2.
+They also present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ with regions surrounded by these curves in ${\mathbb{R}}^{2}$.
+Remark 4 is on ”STEP 1 in our proof of Main Theorems 1 and 2”.
 This is also independent of our present study.
-Here we also add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33].
+Here we also give comments on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33].
 See also [42].
@@
 We do not present elementary notions on graphs precisely.
-We except related knowledge.
-We also note that we omit several rigorous discussions in Remark 3 and that for our main ingredients of the present paper we do not need to understand this rigorously.
-However, we give comments that this is also related to Remark 1 and 2.
+We except related elementary knowledge.
+We also note that we omit several rigorous discussions in Remark 4 and that for our main ingredients of the present paper, we do not need to understand this rigorously.
+We also give comments that this is also related to Remarks 2 and 3.
 Remark 4.
@@
 An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves.
-The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation is considered.
-Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$.
+The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages of all single points for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation on the closure $\overline{R}$ is defined here.
+The vertex set of the graph can be and is defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$.
 The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
@@
     2. (b)
-      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
-A desired region $R_{K}$ can be also obtained by an approximation of the Weierstrass type in the class $C^{2}$.
-According to arguments similar to ones in ”Special Case 5” of [21], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
-The following is a remark on Main Theorem 3.
+      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of some real polynomial function.
+Our desired region $R_{K}$ can be also obtained by a kind of approximation of the Weierstrass type in the class $C^{2}$.
+According to arguments similar to ones in [21, Special Case 5], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial function.
+The following is on Main Theorem 3 and related to some theory of diffeomorphism groups and differential topolgy of manifolds.
 Remark 5.
-We can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively.
-In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
+In Main Theorem 3, let $m=n$.
+By respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively, $M$ and $M_{0}$ are shown to be diffeomorphic to the so-called double of the manifold $\bar{N_{0}}$.
 The following is closely related to one of central problems in real algebraic geometry. This may be closely related to our future work.
 Remark 6.
-Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces has been difficult and important in real algebraic geometry.
+Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces of a general dimension has been difficult and important in real algebraic geometry.
 This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools.
 At present, even explicit construction and classifications for cases of real algebraic curves are difficult.
-Our pioneering studies of the construction respect unit spheres and canonical projections of them where we do not restrict attention to $1$-dimensional cases.
+Our pioneering studies of the construction respect the unit spheres and the canonical projections of them and we do not restrict attention to the $1$-dimensional case.
 Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.
@@
 However, to study more, we need to understand the theory more.
-We close our paper by a short remark on [18].
-Remark 6.
-[18] has no essential errors and we have also found that we need several more rigorous exposition.
-Some of these important arguments are discussed in the present paper.
-We will revise the version and submit our revised version to the refereed journal, based on the presented comments by editors and referees on [18] and our new arguments and idea in the present paper.
-Note also that the content of [18] is independent of the content of the present paper essentially.
-We have also found we can revise [18] in a self-contained way although we need to respect some of our paper.
-We will do so as we can.
 4. Acknowledgement.